### v1

# Feature Coding in the Era of Large Models: Dataset, Test Conditions, and Benchmark

Abstract
Large models have achieved remarkable performance across various tasks, yet they incur significant computational costs and privacy concerns during both training and inference. Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers. To optimize information exchange, feature coding methods are applied to reduce transmission and storage overhead. Despite its importance, feature coding for large models remains an under-explored area. In this paper, we draw attention to large model feature coding and make three contributions to this field. First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models. Second, we establish unified test conditions, enabling standardized evaluation pipelines and fair comparisons across future feature coding studies. Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset. These contributions aim to advance the field of feature coding, facilitating more efficient large model deployment. All source code and the dataset will be made available on GitHub.
Figure 1: Three application scenarios for feature coding in large model deployment. Cloud-centralized training: Features generated by foundation models are encoded and stored in a feature database. This database supports various downstream tasks, reducing computational demands by avoiding repeated inference of large foundation models. Cloud-edge distributed training: To protect user data privacy, raw data is first processed by an edge model to generate features, which are then encoded and transmitted to the cloud. The cloud decodes these features to complete the remaining forward propagation. Cloud-edge distributed inference: A large model is partitioned across multiple parts, with portions offloaded to edge devices to distribute the computational load. The transmitted features ensure privacy protection and enable customized downstream tasks. Please refer to Sec. 3 for detailed illustrations.
1 Introduction
Large models have revolutionized artificial intelligence (AI) with their impressive performance. However, they require substantial training data and extensive computational resources, posing significant challenges for practical deployment [18]. To address these challenges, distributed deployment has been proposed as a promising solution [32, 37, 43, 46, 22, 14]. By distributing the computational workload across multiple platforms, this approach reduces resource strain on individual systems and enables the integration of private client data into training while maintaining data security [43, 4]. As model scale continues to grow, distributed deployment is expected to become a mainstream strategy for large model applications.
In distributed deployment, effective information exchange between model segments is essential. During forward propagation, features act as information carriers, passing features generated by early segments to later segments to complete inference. Given the vast amount of data in large model applications, encoding features to a smaller size to reduce storage and transmission burden is necessary. Analogous to image and video coding, feature coding has been proposed to address this problem [6, 17, 16, 38].
Existing feature coding research has three key limitations: narrow model types, limited task diversity, and constrained data modalities. Most existing feature coding efforts focus on small-scale, convolution-based networks, such as ResNet, ignoring large models. Although discriminative models, such as DINOv2 [28] and SAM [20], represent a significant branch of large models, they are still underrepresented in feature coding research. Generative models also have received little attention in this area. Furthermore, existing research primarily targets visual features, with limited integration of features from large language models (LLMs) based on textual inputs. This gap indicates a need for broader exploration across both model and data modalities to advance feature coding.
In this work, we shift feature coding research from small-scale to large-scale models. To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity. Using this dataset, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
To establish a foundation for large model feature coding, we propose two baseline methods and evaluate them on our dataset. The baselines leverage two primary image coding techniques: the handcrafted VVC Intra coding [2] and the learned Hyperprior method [1]. For each method, we introduce pre-processing and post-processing modules tailored to feature coding. Our analysis provides insights into baseline performance and highlights promising research directions.
In summary, our main contributions are as follows:
  • We emphasize the importance of feature coding in large model deployment, identifying relevant application scenarios and extending feature coding research from small- to large-scale models.
  • We construct a diverse feature dataset encompassing three model types, five tasks, and two data modalities, along with unified test conditions to support future feature coding research.
  • We propose and benchmark two feature coding methods on the dataset, offering baseline evaluations and insights for future research in this field.
Model Parameters Task Split Point Feature Shape Source Data
DINOv2 1.1B Image Classification $40^{th}$ ViT Block $257\times 1536$ ImageNet, 100 samples
Semantic Segmentation $40^{th}$ ViT Block $2\times 1370\times 1536$ VOC2012, 20 samples
Depth Estimation $10^{th}$, $20^{th}$, $30^{th}$, $40^{th}$ ViT Blocks $2\times 4\times 1611\times 1536$ NYU-Depth-v2, 16 samples
Llama3 8B Common Sense Reasoning $32^{th}$ Decoder Layer $N\times 4096$ Arc-Challenge, 100 samples
SD3 8B Text-to-Image Synthesis Input Layer of VAE Decoder $16\times 128\times 128$ COCO2017, 100 samples
Table 1: Summary of the model and task selection, split point decision, and source data collection.
2 Background and motivation
2.1 Feature coding
Visual signal coding encodes and reconstructs the original visual data [33, 44, 15, 25, 34]. Unlike visual signal coding, feature coding refers to the process of encoding features into bitstreams. Originally introduced within the Coding for Machines framework [42, 10], feature coding is one of two primary branches alongside visual coding. One step in time ahead of feature coding, visual signal coding has already been explored in the context of large models [35].
Despite recent progress, feature coding research currently faces three limitations. First, there is a restricted variety of model types in use. The existing feature extraction relies on small-scale, CNN-based models such as ResNet series [23, 31, 19, 24, 3]. However, as Transformer architectures now dominate large model research [36, 11, 12, 28], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide. Second, the field is constrained by the limited scope of task types. Existing studies mainly address discriminative tasks like image classification and segmentation [7, 13, 45, 41, 5, 27], while generative tasks remain largely unexplored. Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential. The third limitation is the constrained range of source modalities. Current feature coding research predominantly focuses on features from visual data, overlooking the increasingly important features derived from textual data, particularly in light of advancements in LLMs. Expanding to multiple modalities is crucial as the diversity of AI applications grows.
Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions. Without a standardized dataset and evaluation conditions, fair comparisons across methods are not feasible. A public dataset ensures consistent input data across studies, while unified test conditions establish a consistent benchmark for evaluating performance.
These limitations motivate us to develop a diverse test dataset and formulate unified test conditions to advance feature coding research.
2.2 Large model deployment
The impressive performance of large models is driven by scaling laws [29, 30], which emphasize the importance of vast model sizes and extensive training data. For example, GPT-3 was built with 175 billion parameters and trained on 499 trillion tokens, both contributing to substantial computational demands. To alleviate these resource requirements, distributed training and inference methods have been proposed in [43, 46, 14].
User privacy presents another critical challenge in client services [4, 26, 40]. Split learning has been introduced as a solution, effectively protecting training data privacy by dividing computation between clients and clouds [43].
In distributed deployment, efficient information exchange between model segments is essential. However, the academic focus has largely overlooked the storage and transmission costs of exchanging these features. Motivated by this gap, our work encodes features into bitstreams to streamline storage and transmission, addressing an essential component of efficient distributed deployment for large models.
3 Feature coding in large models
This section presents three application scenarios of feature coding in large model deployment, illustrated in Fig. 1. Each scenario highlights the role of feature coding in optimizing deployment efficiency.
3.1 Large model training
Large models achieve their high performance through extensive training on large-scale datasets. Due to the limited computational capacity of edge devices, cloud-centralized training is commonly utilized. However, in sensitive domains like healthcare, high-quality public data may be scarce, necessitating access to private data through cloud-edge collaborative training.
3.1.1 Cloud-centralized training
In cloud-centralized training, source databases are maintained in the cloud, where intensive computational demands is the main challenge. Increasingly, complex tasks are addressed through the collaborative use of multiple models. For instance, in vision-language tasks, visual data is initially processed by a vision model, and its features are subsequently passed to an LLM for advanced reasoning. Similarly, foundation models often serve as core backbones for a range of tasks. For example, features extracted from DINOv2 are repurposed for downstream tasks like segmentation [9]. In such collaborative workflows, features generated by one model become input to other models. However, during training, each data sample may undergo repeated inference through the foundation model, leading to a substantial unnecessary computational load.
To mitigate this, an effective solution involves generating features once, storing them, and reusing them across tasks. The workflow is outlined on the left of Fig. 1. First, a foundation model extracts features from the source database, which are then encoded and stored in a feature database. For subsequent task-specific training, only the precomputed feature database is needed, enabling efficient access to features without redundant feature extraction. This significantly reduces computational load and accelerates the training of downstream tasks. However, raw feature data requires substantial storage space, especially with large datasets. To address this, we propose using feature coding to compress raw features into compact bitstreams, reducing storage demands and facilitating efficient, large-scale training.
$SP_{DM1}$
$SP_{DM2}$
$p2$
$p3$
$SP_{DM3}$
$SP_{DM4}$
$p4$
$p5$
$SP_{GS}$
$SP_{H}$
$C5$
$C5$-ReLU
Figure 2: Frequency and CDF comparisons between features in the proposed dataset and commonly used existing features. The proposed features exhibit distinct distributions compared to existing features, highlighting the necessity and value of the dataset. In addition, the proposed features demonstrate greater diversity, enhancing their representativeness and suitability for long-term use. Frequencies are scaled logarithmically for better visualization.
3.1.2 Cloud-edge distributed training
In addition to computational resources, large model training demands access to large-scale datasets, which are often only available to a few giant high-tech companies. Furthermore, sensitive data, such as medical records and financial information, cannot be freely shared due to privacy concerns and regulatory constraints. To leverage distributed, privately owned data, split learning has emerged as a viable solution [37]. Split learning divides the training process into two phases, with the bulk of the computational load handled by the cloud, while edge devices, which are often resource-constrained, handle a smaller portion of the task. This approach minimizes the computational demand on the edge while enabling collaboration with powerful cloud resources.
The cloud-edge distribution training pipeline is depicted in the second diagram of Fig. 1. The whole pipeline is divided into two parts: edge models and cloud-side large models. During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud. Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the large model. In the backward pass, gradients are computed and backpropagated along the reverse path. To optimize this process (reduce bandwidth usage and minimize transmission time), it is critical to apply feature coding before transmission.
3.2 Cloud-edge distributed inference
Once large models are trained, they are deployed in products to provide client services. For a product or service to succeed, two primary concerns must be addressed: supporting a high volume of simultaneous cloud server access and protecting user privacy. For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes. A practical solution is to offload part of this computational demand to edge/client devices. For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec. 3.1.2, can apply.
The cloud-edge distributed inference pipeline is illustrated in the diagram on the right of Fig. 1. In the upload phase, source data is converted into features by an edge model on the client side. These features are then encoded and transmitted to the cloud, where they are further processed by large models. In the download phase, the processed features are sent back to the client to complete a specific task. Depending on the task head used, various functionalities can be performed with the returned features. For example, Stable Diffusion 3 (SD3) features derived from an image can be used to synthesize video or generate segmentation masks. This approach not only reduces the computational load on the cloud server but also enhances data privacy by keeping raw user data on the edge/client side.
$SP_{DM1}$
$SP_{DM2}$
$SP_{DM3}$
$SP_{DM4}$
$SP_{GS}$
$SP_{H}$
$p2$
$p3$
$p4$
$p5$
$C5$
$C5$-ReLU
Figure 3: Visualization of the original feature blocks and their corresponding DCT blocks (absolute values). In the feature domain, the proposed features exhibit higher vertical redundancy, whereas the existing features vary smoothly in both directions. In the DCT domain, the proposed features display more dispersed energy distributions. These distinct redundancy characteristics indicate different coding properties.
4 Dataset construction and analysis
4.1 Dataset construction
To guarantee the dataset’s representativeness and support the long-term study of feature coding research, we curate the test dataset with consideration of three key aspects. Table 1 outlines our selected models and tasks, split points, and source data.
4.1.1 Model and task selection
This study expands the scope of feature coding research by encompassing discriminative, generative, and hybrid models. While prior work has primarily concentrated on discriminative, convolution-based models, the advent of Transformer architectures and generative models calls for a more comprehensive dataset to capture a wider range of tasks and feature types. To meet this need, we select three representative models: DINOv2 [28], Llama3 [11], and SD3 [12]. Llama3 and SD3 represent two core categories within generative models, highlighting their respective strengths in different input-output modalities. Specifically, DINOv2 and Llama3 handle visual and textual inputs, while SD3 processes textual inputs to generate visual outputs. The selected models offer comprehensive feature representations across visual, textual, and cross-modal applications.
For DINOv2, we include the image classification, semantic segmentation, and depth estimation tasks. Llama3 addresses the common sense reasoning task, and SD3 focuses on text-to-image synthesis. This selection covers three visual tasks, one textual task, and one text-to-visual task, forming a versatile foundation for feature coding analysis across diverse task and model types.
4.1.2 Split point decision
This section describes the selection of split points for each model. For DINOv2, we define two types of split points: 1) $SP_{DS}$, the output of the $40^{th}$ ViT layer, and 2) $SP_{DM}$, which aggregates outputs from the $10^{th}$, $20^{th}$, $30^{th}$, and $40^{th}$ ViT layers. $SP_{DS}$ is used for tasks utilizing single-layer features, while $SP_{DM}$ supports tasks that require multi-layer features. In this study, $SP_{DS}$ is applied to image classification and semantic segmentation, and $SP_{DM}$ to depth estimation. The resulting features are designated as $F_{Cls}$, $F_{Seg}$, and $F_{Dpt}$, respectively.
For Llama3, the split point $SP_{G}$ is set at the output of the $32^{nd}$ decoder layer. Features extracted at $SP_{G}$ are versatile, supporting both task-specific heads and integration with other large models [39]. In the common sense reasoning task, $SP_{G}$ produces features of size $N\times 4096$, where $N$ represents the number of tokens.
For SD3, we establish the input layer of the VAE decoder as the split point $SP_{H}$. This split point generates features of size $16\times 128\times 128$, denoted as $F_{TTI}$, which are then passed to the VAE decoder for image synthesis.
4.1.3 Source data collection
Inspired by the video coding standard VVC [2], where test sequences are divided into 8 classes with 3–5 video sequences each, we organize our feature test dataset into five categories. Each category is supported by a curated subset of public datasets used as source data.
For image classification, we select 100 images from ImageNet [8], each representing a unique class accurately classified by DINOv2. For semantic segmentation, we collect 20 images from VOC2012, ensuring coverage of all 20 object classes. For depth estimation, we source 16 images from NYU-Depth-v2, adhering to the train-test split proposed in [21] and selecting one image for each scene category. In the common sense reasoning task, we utilize 100 samples from Arc-Challenge dataset, chosen for their longest input prompts and correct prediction by Llama3. For text-to-image synthesis, we first select 100 images from COCO2017 and then collect their longest captions.
4.2 Feature analysis
In this section, we analyze and compare the proposed features with those commonly used in existing research to clarify the necessity and importance of proposing a new dataset.
4.2.1 Distribution analysis
Frequency distributions and cumulative distribution functions (CDFs) of the proposed and commonly used existing features are visualized in Fig. 2. Features $C5$ and $C5$-ReLU are derived from Pytorch official ResNet50 pre-train on ImageNet, while $P$ features (including $p2$, $p3$, $p4$, and $p5$) are generated by the ResNeXT101-based MaskRCNN pre-trained on COCO2017.
The proposed features exhibit distinct distributions compared to existing features, underscoring the necessity and value of the dataset. For multi-layer split points, four key distinctions between $SP_{DM}$ and $P$ are observed: (1) $SP_{DM}$ distribution range expands progressively with deeper layers; (2) $SP_{DM}$ distribution is more asymmetric; (3) $SP_{DM}$ feature values are more concentrated, as shown in the CDFs; and (4) $SP_{DM}$ features contain more regions with sparse feature points. For single-layer split points, $SP_{GS}$ has a more concentrated distribution, while $SP_{H}$ shows more peaks. Unlike $C5$-ReLU, the proposed dataset omits similar features, as ReLU is infrequently used in large models.
The proposed features demonstrate the high diversity of the dataset, enhancing its representativeness and suitability of long-term research.
Task Head Metric
Cls Norm + Linear1 Accuracy
Seg Norm + Linear1 mIoU
Dpt Linear4 RMSE
CSR Norm + Linear + Softmax Accuracy
TTI VAE Decoder CLIP Score
Table 2: Summary of task head and accuracy metric settings. Specific definitions of Heads can be found in the original papers [28, 11, 12].
4.2.2 Redundancy analysis
We conduct a redundancy analysis on feature blocks, using $7\times 7$ blocks for $C5$ and $C5$-ReLU and $16\times 16$ blocks for other features. We visualize the original feature blocks and their DCT blocks (absolute values) in Fig. 3. In the feature domain, clear spatial redundancy is observed in both $SP_{DM}$ and $SP_{H}$ features, with $SP_{H}$ exhibiting redundancy in both horizontal and vertical directions, while $SP_{DM}$ shows stronger vertical redundancy. By contrast, $SP_{GS}$ features vary rapidly in both directions, resulting in weak spatial redundancy. This is because Llama3 only processes textual data, which lacks inherent spatial correlation. In comparison, existing features exhibit smoother spatial variation and thus high spatial redundancy. We attribute this to the translation invariance of convolutional layers, which preserve the relative positions of pixels in the feature domain, resulting in consistent redundancy patterns. In contrast, transformer-based vision models partition images into patches before transforming them, shifting spatial redundancy primarily to the vertical direction, as seen in $SP_{DM}$.
In the DCT domain, the spatial redundancy is reflected by energy distribution: the proposed features display more distracted energy, whereas existing features concentrate energy in the top-left DC component. Notably, $SP_{H}$ has the highest energy concentration, as it looks similar to images.
The distinct redundancy characteristics indicate different requirements for feature coding methods, which should be explored in future studies.
Figure 4: The pipeline of the proposed baseline feature coding method. The quantization and de-quantization modules are only used in the VTM baseline.
Task Image Classification Semantic Segmentation Depth Estimation Common Sense Reasoning Text-to-Image Synthesis
Metric BPFE Accuracy MSE BPFE mIoU MSE BPFE RMSE MSE BPFE Accuracy MSE BPFE CLIP Score MSE
Quantization 10 100 2.8974 10 79.93 1.8135 10 0.4941 0.5737 10 100 0.0042 10 30.07 0.00001
VTM Baseline 1.90 100 3.0117 1.78 79.68 1.9279 1.67 0.5341 0.6103 2.69 99 0.0114 1.30 30.04 0.0029
0.98 99 3.2324 0.90 78.91 2.1438 0.79 0.6925 0.6763 1.83 100 0.0267 0.65 29.95 0.0078
0.21 81 3.7424 0.24 73.53 2.6032 0.24 1.0684 0.8033 0.88 98 0.0826 0.26 29.50 0.0171
0.04 18 4.0751 0.04 55.05 3.0304 0.05 1.4053 0.9302 0.16 81 0.1900 0.10 27.43 0.0290
0.01 2 4.4588 0.01 32.64 3.3641 0.01 1.5542 1.0313 0.04 26 0.2483 0.04 24.42 0.0436
Hyperprior Baseline 1.99 92 3.5279 1.72 78.44 2.2636 1.51 0.5691 0.6588 5.50 95 0.0674 1.40 30.20 0.0054
1.11 91 3.7724 1.30 77.85 2.3423 1.01 0.6410 0.6830 2.84 89 0.0796 0.62 29.55 0.0122
0.89 86 3.9035 0.54 76.39 2.6461 0.43 0.9442 0.7818 1.46 83 0.1474 0.26 28.21 0.0216
0.37 29 4.3309 0.12 62.69 3.0935 0.08 1.1809 1.0830 1.30 67 0.1337 0.14 26.68 0.0294
0.23 11 4.8063 0.03 40.92 3.6286 0.01 2.5775 1.1876 1.19 32 0.1582 0.07 24.27 0.0404
Table 3: Rate-accuracy performance evaluations on the two baseline methods (MSE is calculated between $F_{o}$ and $\hat{F}_{r}$).
5 Unified test conditions
Even with a test dataset, fair comparisons are still challenging without identical test conditions. To address this, we establish unified test conditions here.
5.1 Bitrate computation
We introduce a new bitrate measurement, Bits Per Feature Point (BPFP), replacing the commonly used Bits Per Pixel (BPP). BPFP is calculated by dividing the total coding bits by the number of feature points.
Our rationale for shifting from BPP to BPFP is twofold. First, BPP is unsuitable for features extracted from non-visual data where pixels do not exist. Second, we believe bitrate should reflect the actual encoded data (the feature itself) rather than the source data. Using BPP to calculate bitrate on source data introduces ambiguity and limits fair comparisons between coding methods. For instance, applying the same BPP to different features extracted from the same image can be misleading. In addition, BPP may incentivize downscaling source images to reduce feature size, yielding lower bitrates that do not represent true coding efficiency. In contrast, BPFP measures bitrate on the feature itself, enabling fair comparisons across coding methods as long as the same feature is used.
5.2 Task accuracy evaluation
To isolate the impact of task heads on task accuracy, we fix their parameters during evaluation. All task head weights are loaded from pre-trained models. The task heads and accuracy metrics are presented in Table 2.
Task Split Point Org. Region Trun. R. (VTM) Trun. R. (Hyperprior)
Cls $SP_{DS}$ [-552.45, 104.18] [-20, 20] [-5, 5]
Seg $SP_{DS}$ [-530.98, 103.22] [-20, 20] [-5, 5]
Dpt $SP_{DM1}$ [-2.42, 3.28] [-1, 1] [-1, 1]
$SP_{DM2}$ [-26.89, 5.03] [-2, 2] [-2, 2]
$SP_{DM3}$ [-323.30, 25.05] [-10, 10] [-10, 10]
$SP_{DM4}$ [-504.43, 102.03] [-20, 20] [-10, 10]
CSR $SP_{G}$ [-78.00, 47.75] [-5, 5] [-5, 5]
TTI $SP_{H}$ [-4.09, 3.05] [-4.09, 3.05] [-4.09, 3.05]
Table 4: Summary of the original and truncation region settings.
6 The proposed baselines and benchmark
6.1 The baseline methods
The pipeline for the proposed feature coding methods is illustrated in Fig. 4. First, the original feature $F_{o}$ is truncated to a smaller range. The quantization operation is performed accordingly. Next, the quantized $F_{q}$ is packed into a 2D feature $F_{p}$. The encoder takes $F_{p}$ as input and outputs decoded feature $\hat{F}_{p}$, which is then unpacked to $\hat{F}_{q}$. $\hat{F}_{q}$ is converted back into a floating-point feature $\hat{F}_{r}$. We propose two baselines: a traditional coding method based on VTM and a learned coding method using the Hyperprior approach.
6.1.1 Pre-processing and post-processing
We propose two pre-processing operations: truncation and quantization. The truncation operation is designed to remove feature points that deviate significantly from the central region. The specific truncated regions for each task are listed in Table 4. Please note that different truncations are used for the two baselines for their distinct coding strategies. The quantization operation uniformly quantizes the truncated features to 10-bit integers. The post-processing only performs de-quantization, which is an inverse process of quantization.
Figure 5: Rate-accuracy-drop comparisons among different tasks. Left: VTM baseline. Right: Hyperprior baseline.
6.1.2 Packing and unpacking
Given the input requirement of VTM, we pack the features into a 2D YUV-400 format. For the Hyperprior encoder, we take the same packing to align with the VTM encoder. $F_{Seg}$, $F_{Dpt}$, and $F_{TTI}$ are packed into the shapes of $2740\times 1536$, $3222\times 6144$, and $512\times 512$, respectively. Please refer to supplementary materials for the packing details.
6.1.3 Encoding and decoding
For the VTM baseline, we employ the VVC reference software VTM-23.3 for feature coding. The Intra coding is applied using the main configuration file encoder_intra_vtm.cfg. The InputChromaFormat is set to 400, and ConformanceWindowMode is enabled. The InternalBitDepth, InputBitDepth, and OutputBitDepth are all set to 10. All other configurations are set to default. In our experiments, five QPs (22, 27, 32, 37, 42) are used.
For the Hyperprior baseline, we modify the input and output to accommodate the 2D features. In addition, we normalized all truncated features into a consistent region of [0,1] before training. Please find detailed training details in the supplementary.
6.2 Rate-accuracy analysis
The rate-accuracy evaluation results are presented in Table 3. We also provide the rate-accuracy curves in the supplementary materials. To further assess the effect of bitrate on accuracy, we introduce a new metric, accuracy drop, defined as the percentage decrease in accuracy. For depth estimation, the reciprocal of RMSE is employed. The bitrate-accuracy-drop curves, shown in Fig. 5, reveal that most tasks exhibit an inflection point where accuracy starts to decrease significantly. However, for depth estimation, this point is observed only in the Hyperprior baseline. In the Hyperprior baseline, inflection points are more distracted across tasks, which may caused by the fact that different models are trained for different tasks. Unlike other tasks, text-to-image synthesis shows smaller accuracy degradation even at low bitrates.
Codec Cls Seg Dpt CSR TTI
VTM Baseline 0.8933 0.8787 0.9904 0.7609 0.9329
Hyperprior Baseline 0.9190 0.9367 0.7741 0.5834 0.9793
Table 5: Linear correlation between MSE and task accuracy.
Configuration Cls Seg Dpt CSR
Truncation Only 100 81.60 0.4992 100
Quantization Only 99 81.37 0.4982 100
Truncation + Quantization 100 79.93 0.4941 100
Quantization + VTM QP22 0 5.35 1.6085 6
Table 6: Ablations on truncation and quantization operations.
6.3 Distortion-accuracy analysis
In this subsection, we examine the linear correlation between feature MSE and task accuracy. We present the coefficients of determination in Table 5. For the VTM baseline, only depth estimation shows a high linear correlation. We believe this arises from its accuracy is measured by RMSE. However, under the Hyperprior baseline, depth estimation does not exhibit a high correlation because of a significant accuracy drop at the lowest bitrate. The common sense reasoning task exhibits an extremely low linear correlation at both baselines, indicating the different characteristics of textual features. This analysis further supports the conclusion that feature MSE is not a good metric for measuring semantic distortion. Developing effective metrics for semantic distortion remains a critical topic in feature coding.
Image Classification Semantic Segmentation
[-5, 5] [-30, 30] [-5, 5] [-30, 30]
BPFE Acc. BPFE Acc. BPFE mIoU BPFE mIoU
2.90 100 0.38 91 2.84 79.83 0.37 76.38
2.05 100 0.06 24 1.98 79.40 0.08 63.46
1.15 100 0.02 14 1.10 78.68 0.01 39.32
Depth Estimation Common Sense Reasoning
[-5, 5], [-5, 5] [-20, 20], [-30, 30] [-2, 2] [-10, 10]
BPFE RMSE BPFE RMSE BPFE Acc. BPFE Acc.
1.42 0.6490 0.58 0.8569 3.16 98 0.84 98
0.73 0.9604 0.17 1.1360 2.30 99 0.15 82
0.33 1.2453 0.04 1.3073 1.44 99 0.04 20
Table 7: Ablation results on truncation regions. $SP_{DM1}$ and $SP_{DM2}$ features are truncated into [-1, 1] and [-2, 2] in all cases, respectively.
6.4 Ablation on truncation and quantization
In this subsection, we examine the impact of truncation and quantization on feature coding. Given the VTM baseline’s stability and deterministic properties, we conduct this ablation study on it specifically. As shown in Table 6, truncation and quantization individually result in little task accuracy loss. When both are applied together, task accuracies experience slight decreases. However, applying only quantization leads to a significant accuracy drop.
Without truncation, feature values are often limited to a small subset within the full range of [0, 1023]. In this scenario, during feature coding, the codec’s internal quantization maps the narrow input range to a single value, introducing significant distortion. However, by applying truncation, this range is expanded, helping to reduce distortion caused by the inner quantization. To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37). As presented in Table 7, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates. Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
7 Conclusion
In this paper, we highlight the critical role of feature coding in deploying large models and make three contributions to advance this field. First, we construct a comprehensive dataset that covers a wide variety of feature types. Second, we establish unified test conditions to support standardized evaluations and ensure fair comparisons for future studies. Third, we introduce two baseline methods, providing foundational benchmarks and insights for researchers exploring this field.
We identify three promising directions for future feature coding research. First, further improvements in coding performance are encouraged. Second, robust methods for measuring semantic distortion in feature coding deserve more attention. Third, developing generic feature coding methods that can handle diverse features across various models is expected.
Supplementary Material
8 Additional dataset analysis
$SP_{DS},F_{Cls}$
$SP_{DS},F_{Seg}$
Figure 6: Visualization of the frequency distribution and CDF for the $F_{Cls}$ and $F_{Seg}$ features.
Proposed Features Existing Features
Source Dataset Split Point Max Min Mean IV GM Split Point Max Min Mean IV GM
ImageNet $SP_{DS}$ 79.01 -502.04 0.0782 17.04 3.97 $C5$-ReLU 7.81 0 0.4594 9839.49 143.22
VOC2012 $SP_{DS}$ 94.88 -485.23 0.0780 13.30 4.30 $C5$ 8.00 -4.24 0.2473 7323.72 131.90
NYUv2 $SP_{DM1}$ 3.25 -2.33 -0.0015 128.91 17.92 $p2$ 10.83 -$1$0.1726 8452.88 159.45
$SP_{DM2}$ 4.99 -26.08 -0.0005 36.96 10.00 $p3$ 8.35 -9.02 -0.1277 9418.53 168.38
$SP_{DM3}$ 23.51 -314.26 0.0042 9.61 2.76 $p4$ 9.06 -9.93 -0.0629 8499.73 158.89
$SP_{DM4}$ 96.04 -493.17 0.0731 12.54 4.35 $p5$ 9.61 -8.22 -0.0295 9402.06 169.13
Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
Captions from COCO2017 $SP_{H}$ 2.19 -1.09 0.5420 22918.26 118.22 / / / / / /
Table 8: Statistics of the proposed features and existing features. (Gradient magnitude values are scaled by $10^{5}$).
We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig. 6 and Fig. 7. These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer. To avoid redundancy in the main paper, these visualizations have been moved to the supplementary material.
Additionally, we compare the proposed features with existing features in terms of statistical characteristics, as presented in Table 8. Specifically, we randomly sample 10 examples from each dataset for statistical analysis, reporting the maximum, minimum, and mean values. In general, the proposed features exhibit broader distribution ranges compared to existing features. For multi-scale features, the distribution range expands with deeper network layers, whereas existing features show smaller variation. In addition, although the proposed features are more asymmetric, their mean values are closer to zero.
To assess spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), with gradients calculated using the Sobel operator. Prior to computation, pixel and feature values are quantized to 10-bit integers. The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range of [7323.72, 9839.49]. A similar trend is observed in gradient magnitude. The larger variations demonstrate the higher diversity of the proposed features. Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features. For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude. This highlights the necessity and importance of including textual features in the proposed dataset.
$SP_{DS},F_{Cls}$
$SP_{DS},F_{Seg}$
Figure 7: Visualization of the original feature blocks and their corresponding DCT blocks for the $F_{Cls}$ and $F_{Seg}$ features.
Task Source Dataset Number of Samples Feature Shape Epoch Learning Rate Lambda
Cls ImageNet 5000 $256\times 256$ 800 1e-4 0.001, 0.0017, 0.003, 0.0035, 0.01
Seg VOC2012 5000 $256\times 256$ 800 1e-4 0.0005, 0.001, 0.003, 0.007, 0.015
Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.045, 10
TTI Captions from COCO2017 50000 $512\times 512$ 60 1e-4 0.005, 0.01, 0.02, 0.05, 0.2
Table 9: Training details of the Hyperprior baseline.
Image Classification
Semantic Segmentation
Depth Estimation
Common Sense Reasoning
Text-to-image Snythesis
$\lambda$-BPFP
Figure 8: Visualization of rate-accuracy performance and $\lambda$-BPFP curves.
9 Baseline methods
9.1 Packing details
9.1.1 Semantic segmentation
In semantic segmentation, an image is pre-processed into multiple $518\times 518$ patches by crop and resize operations. These patches are fed into DINOv2, which produces features of shape $1370\times 1536$. In the proposed dataset, each image is pre-processed into 2 patches, resulting in features ($F_{Seg}$) with a shape of $2\times 1370\times 1536$. We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
9.1.2 Depth estimation
In depth estimation, an image is first horizontally flipped. Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2. For each image, DINOv2 generates $4\times 1611\times 1536$ embedding features at the split point $SP_{DM}$. Thus, the initial shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$, where the two dimensions correspond to the original and flipped images. Packing $F_{Dpt}$ is performed in two steps. First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width. Second, the two $1611\times 6144$ features (from the original and flipped images) are vertically stacked, resulting in a final feature shape of $3222\times 6144$.
9.1.3 Text-to-image synthesis
For text-to-image synthesis, the original feature $F_{TTI}$ has a shape of $16\times 128\times 128$, where 16 represents the number of channels. We divide the 16 channels into 4 subgroups. Within each subgroup, the $128\times 128$ features are horizontally concatenated to form a $128\times 512$ feature. Finally, the four $128\times 512$ features are vertically stacked to produce a final packed feature of size $512\times 512$.
9.2 Training details
9.2.1 Feature Pre-processing
Interestingly, we encountered the storage problem described in the main text during the training of feature coding models. For example, just 20000 original $F_{Dpt}$ features occupy 1474 GB of storage. As the need for additional training data grows, storage demands could escalate significantly. This often exceeds the resources available to researchers in academic institutions.
To address this issue, we crop the features into smaller dimensions to facilitate feature coding training. The specific cropping configurations are detailed in Table 9. However, it is important to note that vertical cropping is not effective for $F_{CSR}$. In our experiments, when $F_{CSR}$ is cropped into $64\times 512$, training becomes unstable, and the coding performance deteriorates significantly. We hypothesize that this limitation arises because $F_{CSR}$ is extracted from an LLM and involves only textual data.
It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training. For large models, it is unreasonable to crop portions of features to accelerate training, as downstream models typically require the entire semantic information. For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation. Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
Additionally, truncation and normalization operations are applied. The truncation regions listed in Table 4 are used. After truncation, all features are normalized into [0, 1].
9.2.2 Loss Function
In our experiments, we train models separately for each task. Across all tasks, the same loss function is employed. The loss function comprises two components: bitrate and feature MSE, defined as:
$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
where $\lambda$ is a scaling factor used to adjust the bitrate. A larger $\lambda$ corresponds to a higher bitrate.
9.2.3 Hyperparameter Setting
The number of training epochs, initial learning rate, and $\lambda$ values are provided in Table 9. The number of training epochs is determined based on the size of the training dataset. In our experiments, all training processes converge to a stable loss. Five different $\lambda$ values are used to achieve various bitrates. The $\lambda$ values are selected to align the bitrates with those achieved by the VTM baseline.
Image Classification
Image Classification
Semantic Segmentation
Semantic Segmentation
Depth Estimation
Depth Estimation
Common Sense Reasoning
Common Sense Reasoning
Text-to-Image Synthesis
Text-to-Image Synthesis
Figure 9: Visualization of the linear relationship between feature MSE and task accuracy. Left: the VTM baseline. Right: the Hyperprior baseline.
Image Classification Semantic Segmentation
BPFE Accu. Accu. Drop BPFE mIoU mIoU Drop
VTM Baseline 1.90 100 0.00% 1.78 79.68 2.35%
0.98 99 1.00% 0.90 78.91 3.30%
0.21 81 19.00% 0.24 73.53 9.89%
0.04 18 82.00% 0.04 55.05 32.54%
0.01 2 98.00% 0.01 32.64 60.00%
Hyperprior Baseline (Trained on Classification) 1.99 92 8.00% 1.53 77.75 4.72%
1.11 91 9.00% 0.88 75.22 7.82%
0.89 86 14.00% 0.70 72.88 10.68%
0.37 29 71.00% 0.24 62.23 23.73%
0.23 11 89.00% 0.09 48.24 40.89%
Hyperprior Baseline (Trained on Segmentation) 2.41 97 3.00% 1.72 78.44 3.87%
1.74 97 3.00% 1.30 77.85 4.59%
0.87 86 14.00% 0.54 76.39 6.39%
0.25 20 80.00% 0.12 62.69 23.17%
0.09 1 99.00% 0.03 40.92 49.85%
Table 10: Rate-accuracy performance comparisons among different feature coding models.
10 Additional experimental results
10.1 Rate-accuracy analysis
We present the rate-accuracy curves in Fig. 8 to give a more straightforward comparison across different tasks. For semantic segmentation and depth estimation, the Hyperprior baseline achieves comparable coding performance to the VTM baseline. However, for the other three tasks, the performance is noticeably poorer. In the case of text-to-image synthesis, this is expected because $F_{TTI}$ shares a distribution similar to that of images and VTM has a higher coding performance in images than the Hyperprior codec. For image classification and common sense reasoning, improved truncation regions or quantization methods could potentially enhance coding performance. However, as feature pre-processing is not the focus of this paper, we leave this exploration to future studies.
For the VTM baseline, we conducted an ablation study on truncation regions. VTM-base, VTM-trun2, and VTM-trun3 correspond to the truncation regions specified in Table 4, the left one of Table 7, and the right one of Table 7, respectively. For depth estimation and common sense reasoning, the tested truncation regions significantly impact coding performance. As with feature pre-processing, the exploration of optimal truncation regions is deferred to future work.
10.2 Distortion-accuracy analysis
We visualize the distortion-accuracy curves in Fig. 9. The VTM baseline performs best in depth estimation, while the Hyperprior baseline achieves the highest linear correlation in text-to-image synthesis. These different behaviors highlight the distinct coding characteristics of the two methods. However, both baselines demonstrate the ineffectiveness of feature MSE as a metric for measuring semantic distortion.
$SP_{DM1},QP=42$
$SP_{DM1},\lambda=0.001$
$SP_{DM2},QP=42$
$SP_{DM2},\lambda=0.001$
$SP_{DM3},QP=42$
$SP_{DM3},\lambda=0.001$
$SP_{DM4},QP=42$
$SP_{DM4},\lambda=0.001$
$SP_{GS},QP=32$
$SP_{GS},\lambda=0.01405$
$SP_{H},QP=42$
$SP_{H},\lambda=0.005$
Figure 10: Frequency and CDF visualization of the original truncated and reconstructed features. Green and red are for the original truncated features. Yellow and black are for the reconstructed features. Left: the VTM baseline. Right: the Hyperprior baseline.
10.3 Exploration of generalization ability
We verify the generalization ability of the two baseline methods. First, we examine the $\lambda$-BPFP relationship and present the results in the bottom-right corner of Fig. 8. The analysis reveals that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties.
To further investigate this, we select the two most similar tasks (image classification and semantic segmentation) based on their $\lambda$-BPFP curves and compare their rate-accuracy performance, as shown in Table 10. For the VTM baseline, although similar bitrates are achieved, the accuracy drop varies significantly at low bitrates. For the Hyperprior baseline, we encode $F_{Cls}$ and $F_{Seg}$ using two groups of models: models trained for image classification and models trained for semantic segmentation. Regardless of which group of models is used, different accuracy drops are observed.
These analyses indicate that the two proposed baselines lack generalization ability across various tasks. The generalization ability of feature coding methods across multiple tasks has received little attention in prior research. By presenting these results, we aim to draw more attention to this important research direction.
10.4 Feature distribution analysis
We visualize the distribution of reconstructed features in Fig. 10, comparing the two baseline methods at similar bitrates. For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range. The reconstructed features appear truncated during the feature coding process. We attribute this truncation to the inner quantization (rounding operation) in the Hyperprior codec, where the transformed latent variables are directly rounded to integers before entropy coding. In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon. A similar truncation phenomenon is observed for $SP_{H}$. For $SP_{GS}$, we visualize features encoded at a higher bitrate and observe that the truncation phenomenon disappears.
This truncation phenomenon highlights the incompatibility between features from large models and learned image coding methods. Two promising research directions emerge: (1) designing a new framework specifically tailored for feature coding, and (2) developing pre-processing techniques to adapt original features for compatibility with image coding methods.

### v3

@@
 Distributed deployment has emerged as a potential solution, but it necessitates the exchange of intermediate information between model segments, with feature representations serving as crucial information carriers.
-To optimize information exchange, feature coding methods are applied to reduce transmission and storage overhead.
+To optimize information exchange, feature coding is required to reduce transmission and storage overhead.
 Despite its importance, feature coding for large models remains an under-explored area.
-In this paper, we draw attention to large model feature coding and make three contributions to this field.
+In this paper, we draw attention to large model feature coding and make three fundamental contributions.
 First, we introduce a comprehensive dataset encompassing diverse features generated by three representative types of large models.
@@
 Third, we introduce two baseline methods derived from widely used image coding techniques and benchmark their performance on the proposed dataset.
-These contributions aim to advance the field of feature coding, facilitating more efficient large model deployment.
-All source code and the dataset will be made available on GitHub.
-Figure 1: Three application scenarios for feature coding in large model deployment.
+These contributions aim to provide a foundation for future research and inspire broader engagement in this field.
+To support a long-term study, All source code and the dataset are made available at https://github.com/chansongoal/FCM-LM/tree/master.
+Figure 1: Three application scenarios for feature coding in large model deployments.
 Cloud-centralized training: Features generated by foundation models are encoded and stored in a feature database.
@@
 Large models have revolutionized artificial intelligence (AI) with their impressive performance.
-However, they require substantial training data and extensive computational resources, posing significant challenges for practical deployment [18].
+However, they require substantial training data and extensive computational resources, posing significant challenges for practical deployments [19].
 To address these challenges, distributed deployment has been proposed as a promising solution [35, 40, 46, 49, 23, 14].
-By distributing the computational workload across multiple platforms, this approach reduces resource strain on individual systems and enables the integration of private client data into training while maintaining data security [43, 4].
-As model scale continues to grow, distributed deployment is expected to become a mainstream strategy for large model applications.
-In distributed deployment, effective information exchange between model segments is essential.
-During forward propagation, features act as information carriers, passing features generated by early segments to later segments to complete inference.
-Given the vast amount of data in large model applications, encoding features to a smaller size to reduce storage and transmission burden is necessary.
+By distributing the computational workload across multiple platforms, this approach reduces the resource strain on individual systems and enables the integration of private client data into training while maintaining data security [46, 4].
+As model scale continues to grow, distributed deployment is expected to become a mainstream strategy for large model deployments.
+In distributed deployments, effective information exchange between model segments is essential.
+During forward propagation, features generated by early segments are passed to later segments to complete the inference.
+Given the vast amount of data in large model applications, it is necessary to encode features to a smaller size to reduce storage and transmission burden.
 Analogous to image and video coding, feature coding has been proposed to address this problem [6, 18, 16, 41].
-Existing feature coding research has three key limitations: narrow model types, limited task diversity, and constrained data modalities.
-Most existing feature coding efforts focus on small-scale, convolution-based networks, such as ResNet, ignoring large models.
-Although discriminative models, such as DINOv2 [28] and SAM [20], represent a significant branch of large models, they are still underrepresented in feature coding research.
-Generative models also have received little attention in this area.
-Furthermore, existing research primarily targets visual features, with limited integration of features from large language models (LLMs) based on textual inputs.
-This gap indicates a need for broader exploration across both model and data modalities to advance feature coding.
-In this work, we shift feature coding research from small-scale to large-scale models.
+Existing feature coding research faces three key limitations: narrow model types, limited task diversity, and constrained data modalities.
+Most existing feature coding efforts focus on small-scale convolution-based networks, such as ResNet, ignoring large models.
+Although existing research emphasizes discriminative models, large models such as DINOv2 [31] and SAM [21] remain underrepresented.
+Generative models also have received little attention in this domain.
+Furthermore, existing research primarily targets visual features and neglects the investigation of textual features from large language models (LLMs).
+This gap indicates a need for broader exploration across both models and data modalities to advance feature coding.
+In this work, we shift the feature coding research from small-scale to large-scale models.
 To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity.
-Using this dataset, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
-To establish a foundation for large model feature coding, we propose two baseline methods and evaluate them on our dataset.
-The baselines leverage two primary image coding techniques: the handcrafted VVC Intra coding [2] and the learned Hyperprior method [1].
-For each method, we introduce pre-processing and post-processing modules tailored to feature coding.
-Our analysis provides insights into baseline performance and highlights promising research directions.
+Furthermore, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
+To lay the groundwork for large model feature coding, we propose two baseline methods and evaluate their applicability to the dataset.
+The baselines leverage two image coding methods: the handcrafted VVC Intra coding [2] and the learning-based Hyperprior method [1].
+We introduce pre-processing and post-processing modules to adapt the baselines to feature characteristics.
+The resulting benchmark and analyses provide valuable insights into large model feature coding and identify promising directions for further exploration.
 In summary, our main contributions are as follows:
-  • We emphasize the importance of feature coding in large model deployment, identifying relevant application scenarios and extending feature coding research from small- to large-scale models.
-  • We construct a diverse feature dataset encompassing three model types, five tasks, and two data modalities, along with unified test conditions to support future feature coding research.
-  • We propose and benchmark two feature coding methods on the dataset, offering baseline evaluations and insights for future research in this field.
+  • We highlight the importance of feature coding in large model deployments, identify relevant application scenarios, and introduce a new research area: large model feature coding.
+  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of data modalities, along with unified test conditions to support long-term feature coding research.
+  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and the reported results suggest promising directions for further research.
 Model Parameters Task Split Point Feature Shape Source Data
-DINOv2 1.1B Image Classification $40^{th}$ ViT Block $257\times 1536$ ImageNet, 100 samples
-Semantic Segmentation $40^{th}$ ViT Block $2\times 1370\times 1536$ VOC2012, 20 samples
-Depth Estimation $10^{th}$, $20^{th}$, $30^{th}$, $40^{th}$ ViT Blocks $2\times 4\times 1611\times 1536$ NYU-Depth-v2, 16 samples
-Llama3 8B Common Sense Reasoning $32^{th}$ Decoder Layer $N\times 4096$ Arc-Challenge, 100 samples
-SD3 8B Text-to-Image Synthesis Input Layer of VAE Decoder $16\times 128\times 128$ COCO2017, 100 samples
+DINOv2 1.1B Image Classification (Cls) $40^{th}$ ViT Block $257\times 1536$ ImageNet, 500 samples
+Semantic Segmentation (Seg) $40^{th}$ ViT Block $2\times 1370\times 1536$ VOC2012, 100 samples
+Depth Estimation (Dpt) $10^{th}$, $20^{th}$, $30^{th}$, $40^{th}$ ViT Blocks $2\times 4\times 1611\times 1536$ NYU-Depth-v2, 80 samples
+Llama3 8B Common Sense Reasoning (CSR) $32^{th}$ Decoder Layer $N\times 4096$ Arc-Challenge, 500 samples
+SD3 8B Text-to-Image Synthesis (TTI) Input Layer of VAE Decoder $16\times 128\times 128$ COCO2017, 500 samples
 Table 1: Summary of the model and task selection, split point decision, and source data collection.
@@
 2.1 Feature coding
-Visual signal coding encodes and reconstructs the original visual data [33, 44, 15, 25, 34].
-Unlike visual signal coding, feature coding refers to the process of encoding features into bitstreams.
 Originally introduced within the Coding for Machines framework [45, 10], feature coding is one of two primary branches alongside visual coding.
-One step in time ahead of feature coding, visual signal coding has already been explored in the context of large models [35].
-Despite recent progress, feature coding research currently faces three limitations.
+Unlike visual signal coding [36, 47, 15, 26, 37, 29], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data (e.g., images) into bitstreams.
+While visual coding has already been explored in the context of large models [38], feature coding remains underdeveloped in this domain.
+Despite recent progress, feature coding research currently faces three key limitations.
 First, there is a restricted variety of model types in use.
-The existing feature extraction relies on small-scale, CNN-based models such as ResNet series [23, 31, 19, 24, 3].
+The existing feature extraction relies on small-scale convolution-based models such as ResNet series [24, 34, 20, 25, 3, 28, 17].
 However, as Transformer architectures now dominate large model research [39, 11, 12, 31], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide.
@@
 Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential.
-The third limitation is the constrained range of source modalities.
-Current feature coding research predominantly focuses on features from visual data, overlooking the increasingly important features derived from textual data, particularly in light of advancements in LLMs.
+The third limitation is the constrained range of source modality.
+Current feature coding research predominantly focuses on features extracted from visual data, overlooking the increasingly important features generated from textual data, particularly in light of advancements in LLMs.
 Expanding to multiple modalities is crucial as the diversity of AI applications grows.
 Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions.
-Without a standardized dataset and evaluation conditions, fair comparisons across methods are not feasible.
-A public dataset ensures consistent input data across studies, while unified test conditions establish a consistent benchmark for evaluating performance.
-These limitations motivate us to develop a diverse test dataset and formulate unified test conditions to advance feature coding research.
+Without a standardized dataset and evaluation conditions, fair comparisons across methods can not be conducted.
+A public dataset ensures consistent input data across studies, while unified test conditions establish a standard pipeline to evaluate performance.
+These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to advance feature coding research.
 2.2 Large model deployment
@@
 To alleviate these resource requirements, distributed training and inference methods have been proposed in [46, 49, 14].
-User privacy presents another critical challenge in client services [4, 26, 40].
-Split learning has been introduced as a solution, effectively protecting training data privacy by dividing computation between clients and clouds [43].
-In distributed deployment, efficient information exchange between model segments is essential.
-However, the academic focus has largely overlooked the storage and transmission costs of exchanging these features.
-Motivated by this gap, our work encodes features into bitstreams to streamline storage and transmission, addressing an essential component of efficient distributed deployment for large models.
+User privacy protection presents another challenge in client services [4, 27, 43].
+Split learning has been proposed as an effective solution to protect data privacy by partitioning models between clients and the cloud [46].
+In distributed deployments, efficient information exchange between model segments is crucial.
+However, academic research has largely overlooked the storage and transmission costs associated with feature exchange.
+To address this gap, we propose to encode features into compact bitstreams to improve the efficiency of large model deployments in distributed environments.
 3 Feature coding in large models
-This section presents three application scenarios of feature coding in large model deployment, illustrated in Fig.
-1.
-Each scenario highlights the role of feature coding in optimizing deployment efficiency.
+In Fig. 1, we illustrate three application scenarios of feature coding in large model deployments, highlighting their roles in optimizing the deployment efficiency of large models.
 3.1 Large model training
 Large models achieve their high performance through extensive training on large-scale datasets.
-Due to the limited computational capacity of edge devices, cloud-centralized training is commonly utilized.
+Cloud-centralized training is commonly employed to leverage the vast computational resources and data in the cloud.
 However, in sensitive domains like healthcare, high-quality public data may be scarce, necessitating access to private data through cloud-edge collaborative training.
 3.1.1 Cloud-centralized training
-In cloud-centralized training, source databases are maintained in the cloud, where intensive computational demands is the main challenge.
-Increasingly, complex tasks are addressed through the collaborative use of multiple models.
-For instance, in vision-language tasks, visual data is initially processed by a vision model, and its features are subsequently passed to an LLM for advanced reasoning.
-Similarly, foundation models often serve as core backbones for a range of tasks.
+In cloud-centralized training, source databases are maintained in the cloud, where intensive computational demands become the main challenge.
+Complex tasks are increasingly addressed through the collaborative use of multiple models.
+For instance, in vision-language tasks, visual data is first processed by a vision model, and its features are then passed to an LLM for advanced reasoning.
+Similarly, foundation models often serve as core backbones for various tasks.
 For example, features extracted from DINOv2 are repurposed for downstream tasks like segmentation [9].
-In such collaborative workflows, features generated by one model become input to other models.
-However, during training, each data sample may undergo repeated inference through the foundation model, leading to a substantial unnecessary computational load.
-To mitigate this, an effective solution involves generating features once, storing them, and reusing them across tasks.
+In such collaborative workflows, features generated by one model become inputs for others.
+However, during training, each data sample may undergo repeated inference, resulting in substantial and unnecessary computational costs.
+To address this, an effective solution is to generate features once, store them, and reuse them across tasks.
 The workflow is outlined on the left of Fig.
@@
 For subsequent task-specific training, only the precomputed feature database is needed, enabling efficient access to features without redundant feature extraction.
-This significantly reduces computational load and accelerates the training of downstream tasks.
-However, raw feature data requires substantial storage space, especially with large datasets.
-To address this, we propose using feature coding to compress raw features into compact bitstreams, reducing storage demands and facilitating efficient, large-scale training.
+This approach significantly reduces computational load and accelerates downstream task training.
+However, raw feature data consumes considerable storage space, particularly with large datasets.
+Therefore, encoding raw features into compact bitstreams becomes essential to reduce storage demands and support efficient large model training.
 $SP_{DM1}$
@@
 3.1.2 Cloud-edge distributed training
-In addition to computational resources, large model training demands access to large-scale datasets, which are often only available to a few giant high-tech companies.
+In addition to computational resources, large model training requires access to large-scale datasets, which are often only available to a few major high-tech companies.
 Furthermore, sensitive data, such as medical records and financial information, cannot be freely shared due to privacy concerns and regulatory constraints.
@@
 This approach minimizes the computational demand on the edge while enabling collaboration with powerful cloud resources.
-The cloud-edge distribution training pipeline is depicted in the second diagram of Fig.
+The cloud-edge distribution training pipeline is depicted in the middle of Fig.
 1.
-The whole pipeline is divided into two parts: edge models and cloud-side large models.
+The whole pipeline is divided into two parts: edge models and cloud models.
 During forward propagation, the edge model processes private data to generate features, which are then encoded into bitstreams and transmitted to the cloud.
-Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the large model.
+Upon receiving the bitstreams, the cloud decodes them back into features and feeds them into the cloud model.
 In the backward pass, gradients are computed and backpropagated along the reverse path.
-To optimize this process (reduce bandwidth usage and minimize transmission time), it is critical to apply feature coding before transmission.
+To optimize this process (reduce bandwidth costs and minimize latency), it is critical to apply feature coding before transmission.
 3.2 Cloud-edge distributed inference
@@
 For the first concern, managing computational load is essential, particularly in generative applications that involve repetitive diffusion processes.
-A practical solution is to offload part of this computational demand to edge/client devices.
+A practical solution is to offload part of this computational demand to edge devices.
 For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec.
 3.1.2, can apply.
-The cloud-edge distributed inference pipeline is illustrated in the diagram on the right of Fig.
+The cloud-edge distributed inference pipeline is illustrated on the right of Fig.
 1.
@@
 For example, Stable Diffusion 3 (SD3) features derived from an image can be used to synthesize video or generate segmentation masks.
-This approach not only reduces the computational load on the cloud server but also enhances data privacy by keeping raw user data on the edge/client side.
+This approach not only reduces the computational load on the cloud server but also enhances data privacy by keeping raw user data on the edge side.
 $SP_{DM1}$
@@
 These distinct redundancy characteristics indicate different coding properties.
+We use $7\times 7$ blocks for $C5$ and $C5$-ReLU and $16\times 16$ blocks for other features.
 4 Dataset construction and analysis
@@
 4.1.1 Model and task selection
-This study expands the scope of feature coding research by encompassing discriminative, generative, and hybrid models.
-While prior work has primarily concentrated on discriminative, convolution-based models, the advent of Transformer architectures and generative models calls for a more comprehensive dataset to capture a wider range of tasks and feature types.
-To meet this need, we select three representative models: DINOv2 [28], Llama3 [11], and SD3 [12].
-Llama3 and SD3 represent two core categories within generative models, highlighting their respective strengths in different input-output modalities.
-Specifically, DINOv2 and Llama3 handle visual and textual inputs, while SD3 processes textual inputs to generate visual outputs.
-The selected models offer comprehensive feature representations across visual, textual, and cross-modal applications.
-For DINOv2, we include the image classification, semantic segmentation, and depth estimation tasks.
-Llama3 addresses the common sense reasoning task, and SD3 focuses on text-to-image synthesis.
-This selection covers three visual tasks, one textual task, and one text-to-visual task, forming a versatile foundation for feature coding analysis across diverse task and model types.
+Given the vast number of large models available, it is impractical to include all of them in a single research paper.
+Therefore, we focus on selecting representative models.
+To ensure the dataset’s representativeness and long-term relevance, we choose one widely recognized model for each type of large model.
+Specifically, we select DINOv2 [31], Llama3 [11], and SD3 [12] as representatives of discriminative models, generative models, and hybrid models, respectively.
+The trend of adopting Transformer architectures in large models is expected to persist in the coming years.
+As a result, the selected models are anticipated to retain their research value and practical utility in the near future.
+DINOv2 and Llama3 handle visual and textual inputs, respectively, while SD3 processes textual inputs to generate visual outputs.
+Together, these models provide comprehensive feature representations across visual, textual, and cross-modal applications.
+For DINOv2, we include image classification (Cls), semantic segmentation (Seg), and depth estimation (Dpt) tasks.
+Llama3 is applied to the common sense reasoning task (CSR), and SD3 is used for text-to-image synthesis (TTI).
+The selection covers three visual tasks, one textual task, and one text-to-visual task, forming a comprehensive foundation for analyzing feature coding across diverse tasks and model types.
 4.1.2 Split point decision
-This section describes the selection of split points for each model.
+The split points are decided to support various downstream tasks.
 For DINOv2, we define two types of split points: 1) $SP_{DS}$, the output of the $40^{th}$ ViT layer, and 2) $SP_{DM}$, which aggregates outputs from the $10^{th}$, $20^{th}$, $30^{th}$, and $40^{th}$ ViT layers.
 $SP_{DS}$ is used for tasks utilizing single-layer features, while $SP_{DM}$ supports tasks that require multi-layer features.
-In this study, $SP_{DS}$ is applied to image classification and semantic segmentation, and $SP_{DM}$ to depth estimation.
-The resulting features are designated as $F_{Cls}$, $F_{Seg}$, and $F_{Dpt}$, respectively.
+In this study, $SP_{DS}$ is applied to Cls and Seg, and $SP_{DM}$ to Dpt.
+The corresponding features are designated as $F_{Cls}$, $F_{Seg}$, and $F_{Dpt}$, respectively.
 For Llama3, the split point $SP_{G}$ is set at the output of the $32^{nd}$ decoder layer.
-Features extracted at $SP_{G}$ are versatile, supporting both task-specific heads and integration with other large models [39].
-In the common sense reasoning task, $SP_{G}$ produces features of size $N\times 4096$, where $N$ represents the number of tokens.
-For SD3, we establish the input layer of the VAE decoder as the split point $SP_{H}$.
+Features extracted at $SP_{G}$ can support both task-specific heads and integration with other large models [42].
+In CSR, $SP_{G}$ produces $N\times 4096$ features, where $N$ is the number of tokens.
+For SD3, we take the input layer of the VAE decoder as the split point $SP_{H}$.
 This split point generates features of size $16\times 128\times 128$, denoted as $F_{TTI}$, which are then passed to the VAE decoder for image synthesis.
 4.1.3 Source data collection
-Inspired by the video coding standard VVC [2], where test sequences are divided into 8 classes with 3–5 video sequences each, we organize our feature test dataset into five categories.
-Each category is supported by a curated subset of public datasets used as source data.
-For image classification, we select 100 images from ImageNet [8], each representing a unique class accurately classified by DINOv2.
-For semantic segmentation, we collect 20 images from VOC2012, ensuring coverage of all 20 object classes.
-For depth estimation, we source 16 images from NYU-Depth-v2, adhering to the train-test split proposed in [21] and selecting one image for each scene category.
-In the common sense reasoning task, we utilize 100 samples from Arc-Challenge dataset, chosen for their longest input prompts and correct prediction by Llama3.
-For text-to-image synthesis, we first select 100 images from COCO2017 and then collect their longest captions.
+We organize our feature test dataset into five classes, each supported by a curated subset of publicly available datasets used as source data.
+For Cls, we select 500 images from ImageNet [8], each representing a unique class accurately classified by DINOv2.
+For Seg, we collect 100 images from VOC2012, ensuring coverage of all 20 object classes.
+For Dpt, we source 80 images from NYU-Depth-v2, adhering to the train-test split proposed in [22] and selecting 3-8 images for each scene category.
+For CSR, we utilize 500 samples from the Arc-Challenge dataset, chosen for their longest input prompts and correct prediction by Llama3.
+For TTI, we first select 500 images from COCO2017 and then collect their longest captions.
+We noticed that learning-based methods may require a dedicated training set.
+We will provide the source code for generating features from any source data to serve as training data.
+Please find out our training data in the supplementary.
 4.2 Feature analysis
-In this section, we analyze and compare the proposed features with those commonly used in existing research to clarify the necessity and importance of proposing a new dataset.
+In this section, we compare the proposed features with those commonly used in existing research to demonstrate the necessity and importance of introducing a new dataset.
 4.2.1 Distribution analysis
-Frequency distributions and cumulative distribution functions (CDFs) of the proposed and commonly used existing features are visualized in Fig.
+We first compare frequency distributions and cumulative distribution functions (CDFs) of the proposed and commonly used existing features in Fig.
 2.
-Features $C5$ and $C5$-ReLU are derived from Pytorch official ResNet50 pre-train on ImageNet, while $P$ features (including $p2$, $p3$, $p4$, and $p5$) are generated by the ResNeXT101-based MaskRCNN pre-trained on COCO2017.
-The proposed features exhibit distinct distributions compared to existing features, underscoring the necessity and value of the dataset.
-For multi-layer split points, four key distinctions between $SP_{DM}$ and $P$ are observed: (1) $SP_{DM}$ distribution range expands progressively with deeper layers; (2) $SP_{DM}$ distribution is more asymmetric; (3) $SP_{DM}$ feature values are more concentrated, as shown in the CDFs; and (4) $SP_{DM}$ features contain more regions with sparse feature points.
+$C5$ and $C5$-ReLU are derived from the official ResNet50 pre-trained on ImageNet, while $P$-layer features are generated by the ResNeXT101-based MaskRCNN pre-trained on COCO2017.
+For multi-layer split points, we observe four key distinctions between $SP_{DM}$ and $P$-layer features: (1) $SP_{DM}$ distribution range expands progressively with deeper layers; (2) $SP_{DM}$ distribution is more asymmetric; (3) $SP_{DM}$ feature values are more concentrated, as evidenced by CDFs; and (4) $SP_{DM}$ features contain more regions with sparse feature points.
 For single-layer split points, $SP_{GS}$ has a more concentrated distribution, while $SP_{H}$ shows more peaks.
-Unlike $C5$-ReLU, the proposed dataset omits similar features, as ReLU is infrequently used in large models.
-The proposed features demonstrate the high diversity of the dataset, enhancing its representativeness and suitability of long-term research.
+The proposed dataset omits $C5$-ReLU-like features, as ReLU is rarely used in large models.
+The proposed features exhibit different distributions compared to the existing features, highlighting the necessity of introducing the new dataset.
+Additionally, the high diversity of the proposed features enhances the dataset’s representativeness and suitability for long-term research.
 Task Head Metric
@@
 CSR Norm + Linear + Softmax Accuracy
-TTI VAE Decoder CLIP Score
-Table 2: Summary of task head and accuracy metric settings. Specific definitions of Heads can be found in the original papers [28, 11, 12].
+TTI VAE Decoder FID
+Table 2: Summary of task head and accuracy metric settings. Find definitions of Heads in the original papers [31, 11, 12].
 4.2.2 Redundancy analysis
-We conduct a redundancy analysis on feature blocks, using $7\times 7$ blocks for $C5$ and $C5$-ReLU and $16\times 16$ blocks for other features.
+Since coding aims to remove redundancy in inputs, we conduct a redundancy comparison on feature blocks.
 We visualize the original feature blocks and their DCT blocks (absolute values) in Fig.
 3.
-In the feature domain, clear spatial redundancy is observed in both $SP_{DM}$ and $SP_{H}$ features, with $SP_{H}$ exhibiting redundancy in both horizontal and vertical directions, while $SP_{DM}$ shows stronger vertical redundancy.
-By contrast, $SP_{GS}$ features vary rapidly in both directions, resulting in weak spatial redundancy.
+In the feature domain, clear spatial redundancy is observed in both $SP_{DM}$ and $SP_{H}$ features.
+$SP_{H}$ exhibits redundancy in both horizontal and vertical directions, while $SP_{DM}$ shows stronger vertical redundancy.
+In contrast, $SP_{GS}$ features vary rapidly in both directions, resulting in weak spatial redundancy.
 This is because Llama3 only processes textual data, which lacks inherent spatial correlation.
-In comparison, existing features exhibit smoother spatial variation and thus high spatial redundancy.
+Existing features, on the other hand, exhibit smoother spatial variation and higher spatial redundancy.
 We attribute this to the translation invariance of convolutional layers, which preserve the relative positions of pixels in the feature domain, resulting in consistent redundancy patterns.
 In contrast, transformer-based vision models partition images into patches before transforming them, shifting spatial redundancy primarily to the vertical direction, as seen in $SP_{DM}$.
-In the DCT domain, the spatial redundancy is reflected by energy distribution: the proposed features display more distracted energy, whereas existing features concentrate energy in the top-left DC component.
-Notably, $SP_{H}$ has the highest energy concentration, as it looks similar to images.
-The distinct redundancy characteristics indicate different requirements for feature coding methods, which should be explored in future studies.
-Figure 4: The pipeline of the proposed baseline feature coding method. The quantization and de-quantization modules are only used in the VTM baseline.
+In the DCT domain, spatial redundancy is reflected in the energy distribution.
+The proposed features display more dispersed energy, whereas existing features concentrate energy in the top-left DC component.
+$SP_{H}$ has the highest energy concentration as it closely resembles images.
+These distinct redundancy characteristics further demonstrate the necessity of the proposed dataset and suggest that future studies should explore feature coding methods tailored to these unique properties.
+Figure 4: The pipeline of the proposed baseline methods.
 Task Image Classification Semantic Segmentation Depth Estimation Common Sense Reasoning Text-to-Image Synthesis
-Metric BPFE Accuracy MSE BPFE mIoU MSE BPFE RMSE MSE BPFE Accuracy MSE BPFE CLIP Score MSE
-Quantization 10 100 2.8974 10 79.93 1.8135 10 0.4941 0.5737 10 100 0.0042 10 30.07 0.00001
-VTM Baseline 1.90 100 3.0117 1.78 79.68 1.9279 1.67 0.5341 0.6103 2.69 99 0.0114 1.30 30.04 0.0029
-0.98 99 3.2324 0.90 78.91 2.1438 0.79 0.6925 0.6763 1.83 100 0.0267 0.65 29.95 0.0078
-0.21 81 3.7424 0.24 73.53 2.6032 0.24 1.0684 0.8033 0.88 98 0.0826 0.26 29.50 0.0171
-0.04 18 4.0751 0.04 55.05 3.0304 0.05 1.4053 0.9302 0.16 81 0.1900 0.10 27.43 0.0290
-0.01 2 4.4588 0.01 32.64 3.3641 0.01 1.5542 1.0313 0.04 26 0.2483 0.04 24.42 0.0436
-Hyperprior Baseline 1.99 92 3.5279 1.72 78.44 2.2636 1.51 0.5691 0.6588 5.50 95 0.0674 1.40 30.20 0.0054
-1.11 91 3.7724 1.30 77.85 2.3423 1.01 0.6410 0.6830 2.84 89 0.0796 0.62 29.55 0.0122
-0.89 86 3.9035 0.54 76.39 2.6461 0.43 0.9442 0.7818 1.46 83 0.1474 0.26 28.21 0.0216
-0.37 29 4.3309 0.12 62.69 3.0935 0.08 1.1809 1.0830 1.30 67 0.1337 0.14 26.68 0.0294
-0.23 11 4.8063 0.03 40.92 3.6286 0.01 2.5775 1.1876 1.19 32 0.1582 0.07 24.27 0.0404
-Table 3: Rate-accuracy performance evaluations on the two baseline methods (MSE is calculated between $F_{o}$ and $\hat{F}_{r}$).
+Metric BPFP Accuracy MSE BPFP mIoU MSE BPFP RMSE MSE BPFP Accuracy MSE BPFP FID MSE
+Original 32 100 0 32 83.39 0 32 0.3695 0 32 100 0 32 0 0
+VTM Baseline 1.94 99.80 2.9499 1.76 80.42 1.8668 2.21 0.5021 0.6108 2.69 99.80 0.0109 1.41 6.14 0.0066
+1.03 97.40 3.1693 0.88 79.31 2.0836 1.27 0.6511 0.6826 1.83 100 0.0260 0.74 19.94 0.0184
+0.23 74.60 3.6993 0.23 73.47 2.5440 0.33 0.9530 0.9225 0.89 98.80 0.0810 0.29 58.42 0.0421
+0.04 25.80 4.0651 0.04 56.53 2.9526 0.03 1.4850 1.0631 0.16 82.20 0.1868 0.11 109.23 0.0715
+0.01 7.20 4.4400 0.01 37.42 3.2762 0.003 2.1174 1.1072 0.04 23.80 0.2455 0.05 171.63 0.1072
+Hyperprior Baseline 2.01 93.20 3.4501 1.71 77.96 2.2010 1.51 0.4256 0.6600 6.34 91.40 0.0776 1.44 25.25 0.0138
+1.13 88.80 3.6966 1.30 77.27 2.2795 1.01 0.4965 0.6844 3.60 87.80 0.0808 0.66 52.12 0.0300
+0.91 83.60 3.8327 0.54 74.82 2.5845 0.43 0.6699 0.7837 1.68 82.40 0.1622 0.28 95.03 0.0541
+0.37 29.00 4.2909 0.12 62.58 3.0151 0.08 1.0351 1.0844 1.50 57.80 0.1355 0.15 124.84 0.0734
+0.23 14.60 4.7814 0.03 37.11 3.5449 0.01 2.7302 1.1866 1.35 34.80 0.1624 0.08 170.01 0.1006
+Table 3: Rate-accuracy (R-A) performance evaluations on the two baseline methods (MSE is calculated between $F_{o}$ and $\hat{F}_{r}$).
 5 Unified test conditions
@@
 5.2 Task accuracy evaluation
-To isolate the impact of task heads on task accuracy, we fix their parameters during evaluation.
-All task head weights are loaded from pre-trained models.
+The decoded features undergo task heads to perform accuracy evaluation.
+To isolate the impact of task heads, we choose simple task heads and fix their parameters during evaluation.
 The task heads and accuracy metrics are presented in Table 2.
 Task Split Point Org. Region Trun. R. (VTM) Trun. R. (Hyperprior)
-Cls $SP_{DS}$ [-552.45, 104.18] [-20, 20] [-5, 5]
-Seg $SP_{DS}$ [-530.98, 103.22] [-20, 20] [-5, 5]
-Dpt $SP_{DM1}$ [-2.42, 3.28] [-1, 1] [-1, 1]
-$SP_{DM2}$ [-26.89, 5.03] [-2, 2] [-2, 2]
+Cls $SP_{DS}$ [-542.30, 94.14] [-20, 20] [-5, 5]
+Seg $SP_{DS}$ [-506.97, 105.95] [-20, 20] [-5, 5]
+Dpt $SP_{DM1}$ [-2.38, 3.27] [-1, 1] [-1, 1]
+$SP_{DM2}$ [-26.44, 5.03] [-2, 2] [-2, 2]
 $SP_{DM3}$ [-323.30, 25.05] [-10, 10] [-10, 10]
-$SP_{DM4}$ [-504.43, 102.03] [-20, 20] [-10, 10]
-CSR $SP_{G}$ [-78.00, 47.75] [-5, 5] [-5, 5]
-TTI $SP_{H}$ [-4.09, 3.05] [-4.09, 3.05] [-4.09, 3.05]
+$SP_{DM4}$ [-504.43, 100.27] [-20, 20] [-10, 10]
+CSR $SP_{G}$ [-71.50, 47.75] [-5, 5] [-5, 5]
+TTI $SP_{H}$ [-5.79, 4.46] [-5.79, 4.46] [-5.79, 4.46]
 Table 4: Summary of the original and truncation region settings.
-6 The proposed baselines and benchmark
+6 Baselines and benchmark
+As a starting point, we introduce two baselines and establish a benchmark.
+We adapt these methods to account for the characteristics of large model features.
+Our goal is to examine their applicability to large model feature coding and provide insights for interested researchers, rather than to develop a completely new method.
 6.1 The baseline methods
-The pipeline for the proposed feature coding methods is illustrated in Fig.
-4.
-First, the original feature $F_{o}$ is truncated to a smaller range.
-The quantization operation is performed accordingly.
+The pipeline for the proposed baselines, illustrated in Fig.
+4, includes pre-processing, core codec, and post-processing modules.
+First, the original feature $F_{o}$ is truncated to a smaller range and then quantized into integers.
 Next, the quantized $F_{q}$ is packed into a 2D feature $F_{p}$.
-The encoder takes $F_{p}$ as input and outputs decoded feature $\hat{F}_{p}$, which is then unpacked to $\hat{F}_{q}$.
-$\hat{F}_{q}$ is converted back into a floating-point feature $\hat{F}_{r}$.
-We propose two baselines: a traditional coding method based on VTM and a learned coding method using the Hyperprior approach.
+The codec takes pre-processed $F_{p}$ as the input and outputs decoded feature $\hat{F}_{p}$.
+Two core codecs are used: a handcrafted codec VTM and a learning-based codec Hyperprior.
+In the post-processing module, $\hat{F}_{p}$ is unpacked to $\hat{F}_{q}$ and then converted back into a floating-point feature $\hat{F}_{r}$.
 6.1.1 Pre-processing and post-processing
-We propose two pre-processing operations: truncation and quantization.
-The truncation operation is designed to remove feature points that deviate significantly from the central region.
+The truncation operation is proposed to remove feature points that deviate significantly from the central region.
 The specific truncated regions for each task are listed in Table 4.
-Please note that different truncations are used for the two baselines for their distinct coding strategies.
+Different truncations are used for the two baselines for their distinct coding strategies.
 The quantization operation uniformly quantizes the truncated features to 10-bit integers.
-The post-processing only performs de-quantization, which is an inverse process of quantization.
-Figure 5: Rate-accuracy-drop comparisons among different tasks. Left: VTM baseline. Right: Hyperprior baseline.
-6.1.2 Packing and unpacking
 Given the input requirement of VTM, we pack the features into a 2D YUV-400 format.
-For the Hyperprior encoder, we take the same packing to align with the VTM encoder.
+The same packing applies in the Hyperprior baseline.
 $F_{Seg}$, $F_{Dpt}$, and $F_{TTI}$ are packed into the shapes of $2740\times 1536$, $3222\times 6144$, and $512\times 512$, respectively.
 Please refer to supplementary materials for the packing details.
+The post-processing performs unpacking and de-quantization, which are inverse processes of packing and quantization.
+Figure 5: Rate-accuracy-drop (R-AD) comparisons among different tasks. Left: VTM baseline. Right: Hyperprior baseline.
 6.1.2 Encoding and decoding
-For the VTM baseline, we employ the VVC reference software VTM-23.3 for feature coding.
+VTM baseline: We employ VTM-23.3 as the codec.
 The Intra coding is applied using the main configuration file encoder_intra_vtm.cfg.
@@
 All other configurations are set to default.
-In our experiments, five QPs (22, 27, 32, 37, 42) are used.
-For the Hyperprior baseline, we modify the input and output to accommodate the 2D features.
-In addition, we normalized all truncated features into a consistent region of [0,1] before training.
-Please find detailed training details in the supplementary.
-6.2 Rate-accuracy analysis
-The rate-accuracy evaluation results are presented in Table 3.
-We also provide the rate-accuracy curves in the supplementary materials.
-To further assess the effect of bitrate on accuracy, we introduce a new metric, accuracy drop, defined as the percentage decrease in accuracy.
-For depth estimation, the reciprocal of RMSE is employed.
-The bitrate-accuracy-drop curves, shown in Fig.
+In our experiments, five QPs {22, 27, 32, 37, 42} are used.
+Hyperprior baseline: We modify the channel of the input and output of the original Hyperprior as 1 to accommodate 2D features. The loss function is defined as:
+$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
+where $\lambda$ is a scaling factor used to adjust the bitrate. Please find more details in the supplementary.
+6.2 Rate-accuracy (R-A) analysis
+The R-A evaluation results are presented in Table 3.
+Both baselines can achieve significant bitrate savings, while the VTM baseline outperforms the Hyperprior baseline.
+To further assess the impact of bitrate on accuracy, we introduce a new metric, accuracy drop (AD), defined as the percentage decrease in accuracy.
+For Dpt, the reciprocal of RMSE is employed.
+The R-AD curves, shown in Fig.
 5, reveal that most tasks exhibit an inflection point where accuracy starts to decrease significantly.
-However, for depth estimation, this point is observed only in the Hyperprior baseline.
+However, for Dpt, this point is observed only in the Hyperprior baseline.
 In the Hyperprior baseline, inflection points are more distracted across tasks, which may caused by the fact that different models are trained for different tasks.
-Unlike other tasks, text-to-image synthesis shows smaller accuracy degradation even at low bitrates.
+TTI is excluded as images generated from the original features have an FID of 0.
+6.3 Distortion-accuracy (D-A) analysis
+We evaluate the linear correlation between feature MSE (D) and task accuracy (A) and present coefficients of determination in Table 5.
+Among the baselines, only TTI exhibits high linear correlations.
+This is reasonable since TTI involves generating images, and the baselines employ image codecs.
+In contrast, CSR fails to achieve a high linear correlation, as its features are derived from textual data.
+Overall, the VTM baseline outperforms the Hyperprior baseline, demonstrating its superior adaptability to large model features.
+Our analysis reveals that feature MSE struggles to characterize semantic distortion, particularly for textual data.
+This finding suggests that feature MSE is not an effective metric for measuring semantic distortion.
+6.4 Complexity analysis
+We run the VTM baseline and Hyperprior baseline on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively.
+The VTM baseline exhibits encoding times ranging from a few seconds to several hours, while its decoding time remains within a few seconds.
+In contrast, the Hyperprior baseline exhibits comparable encoding and decoding times, most of which are on the order of milliseconds.
+Detailed results are provided in the supplementary.
+We note that runtime depends heavily on both hardware capabilities and implementation details.
+A well-optimized codec and high-performance hardware can significantly accelerate feature coding.
 Codec Cls Seg Dpt CSR TTI
-VTM Baseline 0.8933 0.8787 0.9904 0.7609 0.9329
-Hyperprior Baseline 0.9190 0.9367 0.7741 0.5834 0.9793
-Table 5: Linear correlation between MSE and task accuracy.
-Configuration Cls Seg Dpt CSR
-Truncation Only 100 81.60 0.4992 100
-Quantization Only 99 81.37 0.4982 100
-Truncation + Quantization 100 79.93 0.4941 100
-Quantization + VTM QP22 0 5.35 1.6085 6
-Table 6: Ablations on truncation and quantization operations.
-6.3 Distortion-accuracy analysis
-In this subsection, we examine the linear correlation between feature MSE and task accuracy.
-We present the coefficients of determination in Table 5.
-For the VTM baseline, only depth estimation shows a high linear correlation.
-We believe this arises from its accuracy is measured by RMSE.
-However, under the Hyperprior baseline, depth estimation does not exhibit a high correlation because of a significant accuracy drop at the lowest bitrate.
-The common sense reasoning task exhibits an extremely low linear correlation at both baselines, indicating the different characteristics of textual features.
-This analysis further supports the conclusion that feature MSE is not a good metric for measuring semantic distortion.
-Developing effective metrics for semantic distortion remains a critical topic in feature coding.
-Image Classification Semantic Segmentation
-[-5, 5] [-30, 30] [-5, 5] [-30, 30]
-BPFE Acc. BPFE Acc. BPFE mIoU BPFE mIoU
-2.90 100 0.38 91 2.84 79.83 0.37 76.38
-2.05 100 0.06 24 1.98 79.40 0.08 63.46
-1.15 100 0.02 14 1.10 78.68 0.01 39.32
-Depth Estimation Common Sense Reasoning
-[-5, 5], [-5, 5] [-20, 20], [-30, 30] [-2, 2] [-10, 10]
-BPFE RMSE BPFE RMSE BPFE Acc. BPFE Acc.
-1.42 0.6490 0.58 0.8569 3.16 98 0.84 98
-0.73 0.9604 0.17 1.1360 2.30 99 0.15 82
-0.33 1.2453 0.04 1.3073 1.44 99 0.04 20
-Table 7: Ablation results on truncation regions. $SP_{DM1}$ and $SP_{DM2}$ features are truncated into [-1, 1] and [-2, 2] in all cases, respectively.
-6.4 Ablation on truncation and quantization
-In this subsection, we examine the impact of truncation and quantization on feature coding.
-Given the VTM baseline’s stability and deterministic properties, we conduct this ablation study on it specifically.
-As shown in Table 6, truncation and quantization individually result in little task accuracy loss.
-When both are applied together, task accuracies experience slight decreases.
-However, applying only quantization leads to a significant accuracy drop.
-Without truncation, feature values are often limited to a small subset within the full range of [0, 1023].
-In this scenario, during feature coding, the codec’s internal quantization maps the narrow input range to a single value, introducing significant distortion.
-However, by applying truncation, this range is expanded, helping to reduce distortion caused by the inner quantization.
-To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37).
-As presented in Table 7, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates.
-Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
+VTM Baseline 0.9413 0.8955 0.8549 0.7490 0.9983
+Hyperprior Baseline 0.9296 0.9280 0.7437 0.4713 0.9996
+Table 5: Distortion-accuracy (D-A) analysis on the two baselines.
+Task Seg CSR TTI
+Models BPFP mIoU BPFP Acc. BPFP FID
+Trained on Seg 1.71 77.96 2.08 83.80 0.97 119.83
+1.30 77.27 1.37 66.60 0.62 208.50
+0.54 74.82 0.78 0.20 0.31 333.75
+0.12 62.58 0.25 0.00 0.12 363.32
+0.03 37.11 0.09 0.00 0.05 354.22
+Trained on TTI 2.58 78.30 2.28 97.60 1.44 25.25
+1.96 76.34 1.46 77.80 0.66 52.12
+1.12 73.37 0.65 32.00 0.28 95.03
+0.53 56.90 0.30 0.00 0.15 124.84
+0.26 42.94 0.11 0.00 0.08 170.01
+Trained on CSR 5.29 78.20 6.34 91.40 4.77 58.19
+3.27 77.94 3.60 87.80 2.59 73.19
+1.80 76.77 1.68 82.40 1.02 196.56
+1.73 75.82 1.50 57.80 1.02 191.69
+1.54 75.80 1.35 34.80 0.94 193.08
+Table 6: Generalizability evaluation on Seg, CSR, and TTI tasks.
+6.5 Generalizability analysis
+We assess the generalizability of trained models by compressing features extracted from one task using models trained on a different task and evaluating their R-A performance.
+We select three tasks and present the results in Table 6.
+Overall, for a given task (along column), three kinds of models behave differently in both bitrate and accuracy.
+For example, models trained on CSR struggle to achieve low bitrates when applied to Seg and TTI, while models trained on Seg fail to achieve low FID in TTI.
+More analyses can be found in the supplementary.
+In practical applications, an effective compression model should be capable of handling diverse inputs.
+Therefore, a compression method with high generalizability is highly desirable.
+6.6 Discussion
+Large model applications require a balance of high accuracy, low complexity, and strong generalizability.
+However, both baselines struggle to maintain high accuracy at low bitrates, underscoring the need for more efficient feature coding strategies tailored to the unique characteristics of large model features.
+In addition, semantic distortion measurement remains a critical challenge in feature coding.
+Advancing semantic distortion metrics could further optimize codec performance and enhance overall coding efficiency.
+The two baseline approaches present distinct advantages and drawbacks.
+The handcrafted approach aligns seamlessly with existing codecs but suffers from excessive runtime, limiting its real-time applicability.
+In contrast, the learning-based approach benefits from efficient GPU acceleration but lacks generalizability, as separate models must be trained for different features.
+Therefore, developing lightweight and generic codecs remains a promising research direction for feature coding.
 7 Conclusion
-In this paper, we highlight the critical role of feature coding in deploying large models and make three contributions to advance this field.
-First, we construct a comprehensive dataset that covers a wide variety of feature types.
-Second, we establish unified test conditions to support standardized evaluations and ensure fair comparisons for future studies.
-Third, we introduce two baseline methods, providing foundational benchmarks and insights for researchers exploring this field.
-We identify three promising directions for future feature coding research.
-First, further improvements in coding performance are encouraged.
-Second, robust methods for measuring semantic distortion in feature coding deserve more attention.
-Third, developing generic feature coding methods that can handle diverse features across various models is expected.
+In this paper, we highlight the importance of feature coding in large model deployments and introduce a new research area: large model feature coding.
+As the first step, we provide a test dataset, unified test conditions, baseline methods, and a benchmark.
+We examine the applicability of the two baselines and identify promising directions for future research.
+Our aim is to encourage collaboration between image/video coding and large model communities to drive the development of feature coding.
 Supplementary Material
-8 Additional dataset analysis
-$SP_{DS},F_{Cls}$
-$SP_{DS},F_{Seg}$
-Figure 6: Visualization of the frequency distribution and CDF for the $F_{Cls}$ and $F_{Seg}$ features.
-Proposed Features Existing Features
-Source Dataset Split Point Max Min Mean IV GM Split Point Max Min Mean IV GM
-ImageNet $SP_{DS}$ 79.01 -502.04 0.0782 17.04 3.97 $C5$-ReLU 7.81 0 0.4594 9839.49 143.22
-VOC2012 $SP_{DS}$ 94.88 -485.23 0.0780 13.30 4.30 $C5$ 8.00 -4.24 0.2473 7323.72 131.90
-NYUv2 $SP_{DM1}$ 3.25 -2.33 -0.0015 128.91 17.92 $p2$ 10.83 -$1$0.1726 8452.88 159.45
-$SP_{DM2}$ 4.99 -26.08 -0.0005 36.96 10.00 $p3$ 8.35 -9.02 -0.1277 9418.53 168.38
-$SP_{DM3}$ 23.51 -314.26 0.0042 9.61 2.76 $p4$ 9.06 -9.93 -0.0629 8499.73 158.89
-$SP_{DM4}$ 96.04 -493.17 0.0731 12.54 4.35 $p5$ 9.61 -8.22 -0.0295 9402.06 169.13
-Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
-Captions from COCO2017 $SP_{H}$ 2.19 -1.09 0.5420 22918.26 118.22 / / / / / /
-Table 8: Statistics of the proposed features and existing features. (Gradient magnitude values are scaled by $10^{5}$).
-We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig.
-6 and Fig.
-7.
-These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer.
-To avoid redundancy in the main paper, these visualizations have been moved to the supplementary material.
-Additionally, we compare the proposed features with existing features in terms of statistical characteristics, as presented in Table 8.
-Specifically, we randomly sample 10 examples from each dataset for statistical analysis, reporting the maximum, minimum, and mean values.
-In general, the proposed features exhibit broader distribution ranges compared to existing features.
-For multi-scale features, the distribution range expands with deeper network layers, whereas existing features show smaller variation.
-In addition, although the proposed features are more asymmetric, their mean values are closer to zero.
-To assess spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), with gradients calculated using the Sobel operator.
-Prior to computation, pixel and feature values are quantized to 10-bit integers.
-The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range of [7323.72, 9839.49].
-A similar trend is observed in gradient magnitude.
-The larger variations demonstrate the higher diversity of the proposed features.
-Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features.
-For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude.
-This highlights the necessity and importance of including textual features in the proposed dataset.
-$SP_{DS},F_{Cls}$
-$SP_{DS},F_{Seg}$
-Figure 7: Visualization of the original feature blocks and their corresponding DCT blocks for the $F_{Cls}$ and $F_{Seg}$ features.
 Task Source Dataset Number of Samples Feature Shape Epoch Learning Rate Lambda
@@
 Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
-CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.045, 10
+CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.15, 10
 TTI Captions from COCO2017 50000 $512\times 512$ 60 1e-4 0.005, 0.01, 0.02, 0.05, 0.2
 Table 7: Training details of the Hyperprior baseline.
-Image Classification
-Semantic Segmentation
-Depth Estimation
-Common Sense Reasoning
-Text-to-image Snythesis
-$\lambda$-BPFP
-Figure 8: Visualization of rate-accuracy performance and $\lambda$-BPFP curves.
-9 Baseline methods
+8 Details of baseline methods
 8.1 Packing details
@@
 These patches are fed into DINOv2, which produces features of shape $1370\times 1536$.
-In the proposed dataset, each image is pre-processed into 2 patches, resulting in features ($F_{Seg}$) with a shape of $2\times 1370\times 1536$.
+In the proposed dataset, each image is pre-processed into 2 patches, resulting in features with a shape of $2\times 1370\times 1536$.
 We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
@@
 Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2.
-For each image, DINOv2 generates $4\times 1611\times 1536$ embedding features at the split point $SP_{DM}$.
-Thus, the initial shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$, where the two dimensions correspond to the original and flipped images.
-Packing $F_{Dpt}$ is performed in two steps.
+For each image, DINOv2 generates $4\times 1611\times 1536$ features at the split point $SP_{DM}$.
+Thus, the shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$.
+$F_{Dpt}$ is packed in two steps.
 First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width.
-Second, the two $1611\times 6144$ features (from the original and flipped images) are vertically stacked, resulting in a final feature shape of $3222\times 6144$.
+Second, the two $1611\times 6144$ features (original first and then flipped images) are vertically stacked, forming a $3222\times 6144$ feature.
 8.1.3 Text-to-image synthesis
@@
 8.2 Training details
-9.2.1 Feature Pre-processing
-Interestingly, we encountered the storage problem described in the main text during the training of feature coding models.
-For example, just 20000 original $F_{Dpt}$ features occupy 1474 GB of storage.
-As the need for additional training data grows, storage demands could escalate significantly.
-This often exceeds the resources available to researchers in academic institutions.
+8.2.1 Training dataset construction
+We construct our training dataset from public datasets.
+The training features are extracted in the same way as the test dataset.
+The source code will be made publicly available, and we encourage researchers to extract features from their own source data.
+As mentioned in the main text, raw features consume a significant amount of storage.
+During training, we encountered storage challenges.
+For example, 20000 original $F_{Dpt}$ features occupy 1474 GB of storage.
+As the demand for additional training data grows, storage requirements could increase substantially, often exceeding the resources available to researchers in academic institutions.
 To address this issue, we crop the features into smaller dimensions to facilitate feature coding training.
 The specific cropping configurations are detailed in Table 7.
-However, it is important to note that vertical cropping is not effective for $F_{CSR}$.
-In our experiments, when $F_{CSR}$ is cropped into $64\times 512$, training becomes unstable, and the coding performance deteriorates significantly.
-We hypothesize that this limitation arises because $F_{CSR}$ is extracted from an LLM and involves only textual data.
 It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training.
-For large models, it is unreasonable to crop portions of features to accelerate training, as downstream models typically require the entire semantic information.
+For large models, cropping features to accelerate training is impractical, as downstream models typically require complete semantic information.
 For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation.
 Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
-Additionally, truncation and normalization operations are applied. The truncation regions listed in Table 4 are used. After truncation, all features are normalized into [0, 1].
-9.2.2 Loss Function
-In our experiments, we train models separately for each task. Across all tasks, the same loss function is employed. The loss function comprises two components: bitrate and feature MSE, defined as:
-$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
-where $\lambda$ is a scaling factor used to adjust the bitrate. A larger $\lambda$ corresponds to a higher bitrate.
+After feature extraction, the same truncation and quantization operations are applied to the training dataset.
+For the Hyperprior baseline, all features are normalized to the range [0, 1] before training.
 8.2.2 Hyperparameter Setting
-The number of training epochs, initial learning rate, and $\lambda$ values are provided in Table 9.
+The hyperparameters, such as the number of training epochs, initial learning rate, and $\lambda$ values, are provided in Table 7.
 The number of training epochs is determined based on the size of the training dataset.
@@
 Five different $\lambda$ values are used to achieve various bitrates.
-The $\lambda$ values are selected to align the bitrates with those achieved by the VTM baseline.
-Image Classification
-Image Classification
+Task Image Classification Semantic Segmentation Depth Estimation Common Sense Reasoning Text-to-Image Synthesis
+Metric QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec.
+VTM Baseline 22 267.20 0.23 22 2301.30 1.11 22 14169.36 5.15 22 336.27 0.24 22 107.98 0.16
+27 275.02 0.22 27 2039.34 0.94 27 10862.76 3.54 27 423.84 0.29 27 87.33 0.15
+32 124.40 0.15 32 1399.71 0.52 32 7776.30 2.33 32 362.98 0.24 32 62.11 0.15
+37 28.00 0.13 37 407.30 0.32 37 1590.27 1.14 37 162.05 0.17 37 32.38 0.24
+42 8.55 0.14 42 70.66 0.27 42 348.68 1.17 42 66.28 0.18 42 15.70 0.15
+Metric $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec.
+Hyperprior Baseline 0.01 0.04 0.06 0.015 0.3 0.38 0.12 1.57 2.05 10 0.07 0.12 0.2 0.03 0.03
+0.0035 0.04 0.05 0.007 0.31 0.4 0.05 1.59 2.05 0.07 0.07 0.09 0.05 0.03 0.03
+0.003 0.04 0.05 0.003 0.31 0.39 0.02 1.55 1.88 0.015 0.06 0.08 0.02 0.03 0.03
+0.0017 0.05 0.06 0.001 0.29 0.35 0.005 1.54 1.84 0.0142 0.07 0.08 0.01 0.03 0.03
+0.001 0.05 0.06 0.0005 0.29 0.36 0.001 1.57 1.91 0.01405 0.06 0.08 0.005 0.03 0.03
+Table 8: Encoding and decoding complexity evaluations. “Enc.” and “Dec.” denote encoding time and decoding time, respectively. The running time is measured in seconds.
+$\lambda$-BPFP
 Semantic Segmentation
-Semantic Segmentation
-Depth Estimation
-Depth Estimation
-Common Sense Reasoning
 Common Sense Reasoning
 Text-to-Image Synthesis
-Text-to-Image Synthesis
-Figure 9: Visualization of the linear relationship between feature MSE and task accuracy. Left: the VTM baseline. Right: the Hyperprior baseline.
-Image Classification Semantic Segmentation
-BPFE Accu. Accu. Drop BPFE mIoU mIoU Drop
-VTM Baseline 1.90 100 0.00% 1.78 79.68 2.35%
-0.98 99 1.00% 0.90 78.91 3.30%
-0.21 81 19.00% 0.24 73.53 9.89%
-0.04 18 82.00% 0.04 55.05 32.54%
-0.01 2 98.00% 0.01 32.64 60.00%
-Hyperprior Baseline (Trained on Classification) 1.99 92 8.00% 1.53 77.75 4.72%
-1.11 91 9.00% 0.88 75.22 7.82%
-0.89 86 14.00% 0.70 72.88 10.68%
-0.37 29 71.00% 0.24 62.23 23.73%
-0.23 11 89.00% 0.09 48.24 40.89%
-Hyperprior Baseline (Trained on Segmentation) 2.41 97 3.00% 1.72 78.44 3.87%
-1.74 97 3.00% 1.30 77.85 4.59%
-0.87 86 14.00% 0.54 76.39 6.39%
-0.25 20 80.00% 0.12 62.69 23.17%
-0.09 1 99.00% 0.03 40.92 49.85%
-Table 10: Rate-accuracy performance comparisons among different feature coding models.
+Figure 6: $\lambda$-BPFP curves and R-A curves of generalizability evaluation on Seg, CSR, and TTI tasks.
 9 Additional experimental results
-10.1 Rate-accuracy analysis
-We present the rate-accuracy curves in Fig.
-8 to give a more straightforward comparison across different tasks.
-For semantic segmentation and depth estimation, the Hyperprior baseline achieves comparable coding performance to the VTM baseline.
-However, for the other three tasks, the performance is noticeably poorer.
-In the case of text-to-image synthesis, this is expected because $F_{TTI}$ shares a distribution similar to that of images and VTM has a higher coding performance in images than the Hyperprior codec.
-For image classification and common sense reasoning, improved truncation regions or quantization methods could potentially enhance coding performance.
-However, as feature pre-processing is not the focus of this paper, we leave this exploration to future studies.
-For the VTM baseline, we conducted an ablation study on truncation regions.
-VTM-base, VTM-trun2, and VTM-trun3 correspond to the truncation regions specified in Table 4, the left one of Table 7, and the right one of Table 7, respectively.
-For depth estimation and common sense reasoning, the tested truncation regions significantly impact coding performance.
-As with feature pre-processing, the exploration of optimal truncation regions is deferred to future work.
-10.2 Distortion-accuracy analysis
-We visualize the distortion-accuracy curves in Fig.
-9.
-The VTM baseline performs best in depth estimation, while the Hyperprior baseline achieves the highest linear correlation in text-to-image synthesis.
-These different behaviors highlight the distinct coding characteristics of the two methods.
-However, both baselines demonstrate the ineffectiveness of feature MSE as a metric for measuring semantic distortion.
 $SP_{DM1},QP=42$
@@
 Right: the Hyperprior baseline.
-10.3 Exploration of generalization ability
-We verify the generalization ability of the two baseline methods.
-First, we examine the $\lambda$-BPFP relationship and present the results in the bottom-right corner of Fig.
-8.
-The analysis reveals that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties.
-To further investigate this, we select the two most similar tasks (image classification and semantic segmentation) based on their $\lambda$-BPFP curves and compare their rate-accuracy performance, as shown in Table 10.
-For the VTM baseline, although similar bitrates are achieved, the accuracy drop varies significantly at low bitrates.
-For the Hyperprior baseline, we encode $F_{Cls}$ and $F_{Seg}$ using two groups of models: models trained for image classification and models trained for semantic segmentation.
-Regardless of which group of models is used, different accuracy drops are observed.
-These analyses indicate that the two proposed baselines lack generalization ability across various tasks.
-The generalization ability of feature coding methods across multiple tasks has received little attention in prior research.
-By presenting these results, we aim to draw more attention to this important research direction.
-10.4 Feature distribution analysis
+9.1 Complexity analysis
+We present the encoding and decoding complexity analyses in Table 8.
+The VTM baseline and Hyperprior baseline are evaluated on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively.
+It is important to note that the VTM baseline is a handcrafted codec and cannot be executed on a GPU.
+For the VTM baseline, encoding time increases as QP decreases and feature resolution increases.
+The block-based VTM codec exhaustively examines multiple partitions and modes during encoding, making it significantly more time-consuming than decoding.
+VTM is one of the reference implementations of the VVC standard.
+In practical deployments, a more efficient VVC codec, such as VVEnc, can be used to accelerate encoding.
+For the Hyperprior baseline, the decoding complexity is slightly higher than the encoding complexity.
+Across different bitrates, the encoding and decoding complexities remain comparable.
+9.2 Generalizability analysis
+Here, we visualize the $\lambda$-BPFP curves and the rate-accuracy curves for the cross evaluation to provide a more intuitive presentation of the results.
+The $\lambda$-BPFP curves reveal that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties.
+From the rate-accuracy curves, we observe that coding models trained for one task struggle to achieve comparable performance when applied to another task.
+The generalizability of feature coding methods across multiple tasks has received limited attention in prior research.
+By presenting these results, we aim to highlight this important research direction and encourage further exploration.
+9.3 Ablation on truncation and quantization
+In this subsection, we examine the impact of truncation and quantization on feature coding.
+As shown in Table 10, truncation and quantization individually result in little task accuracy loss.
+When both are applied together, task accuracies experience slight decreases.
+However, applying the VTM baseline to quantized features leads to a significant reduction in accuracy.
+Without truncation, feature values are often limited to a small subset within the full range of [0, 1023].
+In this scenario, during feature coding, the codec’s internal quantization maps values with a small range to a single value, introducing significant distortion.
+By applying truncation, this range is expanded, which helps mitigate distortion caused by VTM’s internal quantization.
+To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37).
+As presented in Table 11, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates.
+Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
+Proposed Features Existing Features
+Source Dataset Split Point Max Min Mean IV GM Split Point Max Min Mean IV GM
+ImageNet $SP_{DS}$ 79.01 -502.04 0.0782 17.04 3.97 $C5$-ReLU 7.81 0 0.4594 9839.49 143.22
+VOC2012 $SP_{DS}$ 94.88 -485.23 0.0780 13.30 4.30 $C5$ 8.00 -4.24 0.2473 7323.72 131.90
+NYUv2 $SP_{DM1}$ 3.25 -2.33 -0.0015 128.91 17.92 $p2$ 10.83 -$1$0.1726 8452.88 159.45
+$SP_{DM2}$ 4.99 -26.08 -0.0005 36.96 10.00 $p3$ 8.35 -9.02 -0.1277 9418.53 168.38
+$SP_{DM3}$ 23.51 -314.26 0.0042 9.61 2.76 $p4$ 9.06 -9.93 -0.0629 8499.73 158.89
+$SP_{DM4}$ 96.04 -493.17 0.0731 12.54 4.35 $p5$ 9.61 -8.22 -0.0295 9402.06 169.13
+Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
+Captions from COCO2017 $SP_{H}$ 3.67 -2.11 0.5664 21130.12 138.53 / / / / / /
+Table 9: Statistics of the proposed features and existing features. (Gradient magnitude values are scaled by $10^{5}$).
+Configuration Cls Seg Dpt CSR
+Truncation Only 100 81.60 0.4992 100
+Quantization Only 99 81.37 0.4982 100
+Truncation + Quantization 100 79.93 0.4941 100
+Quantization + VTM QP22 0 5.35 1.6085 6
+Table 10: Ablations on truncation and quantization operations.
+9.4 Distribution analysis on decoded features
 We visualize the distribution of reconstructed features in Fig.
-10, comparing the two baseline methods at similar bitrates.
+7, comparing the two baseline methods at similar bitrates.
 For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range.
 The reconstructed features appear truncated during the feature coding process.
-We attribute this truncation to the inner quantization (rounding operation) in the Hyperprior codec, where the transformed latent variables are directly rounded to integers before entropy coding.
+We attribute this truncation to the Hyperprior’s internal quantization (rounding operation), where the transformed latent variables are directly rounded to integers before entropy coding.
 In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon.
 A similar truncation phenomenon is observed for $SP_{H}$.
-For $SP_{GS}$, we visualize features encoded at a higher bitrate and observe that the truncation phenomenon disappears.
-This truncation phenomenon highlights the incompatibility between features from large models and learned image coding methods.
-Two promising research directions emerge: (1) designing a new framework specifically tailored for feature coding, and (2) developing pre-processing techniques to adapt original features for compatibility with image coding methods.
+For $SP_{GS}$, the features are encoded at a higher bitrate and the truncation phenomenon is not observed.
+10 Additional dataset analysis
+In addition to the visualizations provided in the main text, we present further statistical analyses of the proposed dataset.
+10.1 Distribution analysis
+We compare the statistical characteristics of the proposed features with those of existing features by randomly sampling 10 examples from each dataset and reporting their maximum, minimum, and mean values in Table 9.
+Overall, the proposed features exhibit a broader distribution range compared to existing features.
+In particular, for multi-scale features, the distribution range expands in deeper network layers, whereas existing features exhibit smaller variations.
+Moreover, although the proposed features tend to be more asymmetric, their mean values remain closer to zero.
+Image Classification Semantic Segmentation
+[-5, 5] [-30, 30] [-5, 5] [-30, 30]
+BPFE Acc. BPFE Acc. BPFE mIoU BPFE mIoU
+2.90 100 0.38 91 2.84 79.83 0.37 76.38
+2.05 100 0.06 24 1.98 79.40 0.08 63.46
+1.15 100 0.02 14 1.10 78.68 0.01 39.32
+Depth Estimation Common Sense Reasoning
+[-5, 5], [-5, 5] [-20, 20], [-30, 30] [-2, 2] [-10, 10]
+BPFE RMSE BPFE RMSE BPFE Acc. BPFE Acc.
+1.42 0.6490 0.58 0.8569 3.16 98 0.84 98
+0.73 0.9604 0.17 1.1360 2.30 99 0.15 82
+0.33 1.2453 0.04 1.3073 1.44 99 0.04 20
+Table 11: Ablation results on truncation regions. $SP_{DM1}$ and $SP_{DM2}$ features are truncated into [-1, 1] and [-2, 2] in all cases, respectively.
+10.2 Redundancy analysis
+To evaluate spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), where gradients are derived using the Sobel operator.
+Before these computations, both pixel and feature values are quantized to 10-bit integers.
+The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range from 7323.72 to 9839.49.
+A similar trend is observed in gradient magnitude.
+These broader variations underscore the higher diversity of the proposed features.
+Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features.
+For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude.
+This observation highlights the necessity and importance of incorporating textual features in the dataset.
+10.3 Visualization
+We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig.
+8 and Fig.
+9.
+These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer.
+To avoid redundancy in the main paper, these visualizations have been moved here.
+$SP_{DS},F_{Cls}$
+$SP_{DS},F_{Seg}$
+Figure 8: Visualization of the frequency distribution and CDF for the $F_{Cls}$ and $F_{Seg}$ features.
+$SP_{DS},F_{Cls}$
+$SP_{DS},F_{Seg}$
+Figure 9: Visualization of the original feature blocks and their corresponding DCT blocks for the $F_{Cls}$ and $F_{Seg}$ features.

### v4

@@
 These contributions aim to provide a foundation for future research and inspire broader engagement in this field.
-To support a long-term study, All source code and the dataset are made available at https://github.com/chansongoal/FCM-LM/tree/master.
+To support a long-term study, all source code and the dataset are made available at https://github.com/chansongoal/LaMoFC.
+footnotetext: This work was supported by the Ministry of Education of Singapore under Grant T2EP20123-0006.
+We acknowledge the support of GPU cluster built by MCC Lab of Information Science and Technology Institution, USTC.
 Figure 1: Three application scenarios for feature coding in large model deployments.
@@
 Most existing feature coding efforts focus on small-scale convolution-based networks, such as ResNet, ignoring large models.
-Although existing research emphasizes discriminative models, large models such as DINOv2 [31] and SAM [21] remain underrepresented.
-Generative models also have received little attention in this domain.
 Furthermore, existing research primarily targets visual features and neglects the investigation of textual features from large language models (LLMs).
@@
 To address the existing limitations, we construct a comprehensive test dataset with features from five tasks across three large models, covering both visual and textual data for increased modality diversity.
-Furthermore, we establish unified test conditions to enable fair coding performance comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
+Furthermore, we establish unified test conditions to enable fair comparisons, including formulated bitrate computation and task-specific evaluation pipelines.
 To lay the groundwork for large model feature coding, we propose two baseline methods and evaluate their applicability to the dataset.
@@
   • We highlight the importance of feature coding in large model deployments, identify relevant application scenarios, and introduce a new research area: large model feature coding.
-  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of data modalities, along with unified test conditions to support long-term feature coding research.
-  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and the reported results suggest promising directions for further research.
+  • We construct a comprehensive feature dataset encompassing three model types, five tasks, and two kinds of source data modalities, along with unified test conditions to support long-term large model feature coding research.
+  • We propose two baselines and evaluate their usability to the large model feature coding. We build a benchmark and suggest promising directions for further research.
 Model Parameters Task Split Point Feature Shape Source Data
@@
 Originally introduced within the Coding for Machines framework [48, 10], feature coding is one of two primary branches alongside visual coding.
-Unlike visual signal coding [36, 47, 15, 26, 37, 29], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data (e.g., images) into bitstreams.
-While visual coding has already been explored in the context of large models [38], feature coding remains underdeveloped in this domain.
-Despite recent progress, feature coding research currently faces three key limitations.
+Unlike visual coding [39, 50, 15, 28, 40, 31, 35, 26, 25], which encodes and reconstructs the original visual data, feature coding involves encoding features extracted from source data into bitstreams.
+While visual coding has already been explored in the context of large models [41, 20], feature coding remains unexplored in this domain.
+Despite recent progress, feature coding research still faces three key limitations.
 First, there is a restricted variety of model types in use.
-The existing feature extraction relies on small-scale convolution-based models such as ResNet series [24, 34, 20, 25, 3, 28, 17].
-However, as Transformer architectures now dominate large model research [39, 11, 12, 31], it is crucial to investigate feature coding for Transformer-based models to capture the rich representations they provide.
+The existing feature coding focuses on small-scale CNN features such as [24, 37, 21, 27, 3, 30, 17].
+However, as Transformer architectures now dominate large model research [42, 11, 12, 33], it is crucial to investigate feature coding for Transformer-based models.
 Second, the field is constrained by the limited scope of task types.
-Existing studies mainly address discriminative tasks like image classification and segmentation [7, 13, 48, 44, 5, 30], while generative tasks remain largely unexplored.
-Given the significance of generative models in modern AI, extending feature coding to include these tasks, such as image synthesis, is essential.
+Existing studies primarily address discriminative tasks such as image classification and segmentation [7, 13, 51, 47, 5, 32], while generative tasks remain largely unexplored.
+Given the significance of generative models, extending feature coding to include these tasks, such as image synthesis, is essential.
 The third limitation is the constrained range of source modality.
@@
 Expanding to multiple modalities is crucial as the diversity of AI applications grows.
-Additionally, two core challenges hinder progress in feature coding for large models: the lack of a public test dataset and unified test conditions.
+Furthermore, two challenges hinder progress in large model feature coding: the lack of a public test dataset and unified test conditions.
 Without a standardized dataset and evaluation conditions, fair comparisons across methods can not be conducted.
 A public dataset ensures consistent input data across studies, while unified test conditions establish a standard pipeline to evaluate performance.
-These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to advance feature coding research.
+These limitations motivate us to construct a diverse test dataset and formulate unified test conditions to support large model feature coding research.
 2.2 Large model deployment
@@
 This approach minimizes the computational demand on the edge while enabling collaboration with powerful cloud resources.
-The cloud-edge distribution training pipeline is depicted in the middle of Fig.
+The cloud-edge distributed training pipeline is depicted in the middle of Fig.
 1.
@@
 In the backward pass, gradients are computed and backpropagated along the reverse path.
-To optimize this process (reduce bandwidth costs and minimize latency), it is critical to apply feature coding before transmission.
+To reduce bandwidth cost and minimize latency, it is critical to apply feature coding before transmission.
 3.2 Cloud-edge distributed inference
@@
 For the second concern, the same strategy in the cloud-edge distributed training, as discussed in Sec.
-3.1.2, can apply.
+3.1.2, can be applied.
 The cloud-edge distributed inference pipeline is illustrated on the right of Fig.
@@
 For TTI, we first select 500 images from COCO2017 and then collect their longest captions.
-We noticed that learning-based methods may require a dedicated training set.
-We will provide the source code for generating features from any source data to serve as training data.
-Please find out our training data in the supplementary.
+We noticed that learning-based methods require a dedicated training set. Please follow the source code to generate training data and find our training data in the supplementary.
 4.2 Feature analysis
@@
 Original 32 100 0 32 83.39 0 32 0.3695 0 32 100 0 32 0 0
-VTM Baseline 1.94 99.80 2.9499 1.76 80.42 1.8668 2.21 0.5021 0.6108 2.69 99.80 0.0109 1.41 6.14 0.0066
+VTM
+Baseline
+1.94 99.80 2.9499 1.76 80.42 1.8668 2.21 0.5021 0.6108 2.69 99.80 0.0109 1.41 6.14 0.0066
 1.03 97.40 3.1693 0.88 79.31 2.0836 1.27 0.6511 0.6826 1.83 100 0.0260 0.74 19.94 0.0184
@@
 0.01 7.20 4.4400 0.01 37.42 3.2762 0.003 2.1174 1.1072 0.04 23.80 0.2455 0.05 171.63 0.1072
-Hyperprior Baseline 2.01 93.20 3.4501 1.71 77.96 2.2010 1.51 0.4256 0.6600 6.34 91.40 0.0776 1.44 25.25 0.0138
+Hyperprior
+Baseline
+2.01 93.20 3.4501 1.71 77.96 2.2010 1.51 0.4256 0.6600 6.34 91.40 0.0776 1.44 25.25 0.0138
 1.13 88.80 3.6966 1.30 77.27 2.2795 1.01 0.4965 0.6844 3.60 87.80 0.0808 0.66 52.12 0.0300
@@
 6 Baselines and benchmark
-As a starting point, we introduce two baselines and establish a benchmark.
-We adapt these methods to account for the characteristics of large model features.
-Our goal is to examine their applicability to large model feature coding and provide insights for interested researchers, rather than to develop a completely new method.
+As a starting point, we introduce two baselines and establish a benchmark. We aim to examine their applicability to large model feature coding and provide insights for interested researchers.
 6.1 The baseline methods
@@
 Hyperprior baseline: We modify the channel of the input and output of the original Hyperprior as 1 to accommodate 2D features. The loss function is defined as:
-$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
+[…]
 where $\lambda$ is a scaling factor used to adjust the bitrate. Please find more details in the supplementary.
@@
 Our aim is to encourage collaboration between image/video coding and large model communities to drive the development of feature coding.
-Supplementary Material
-Task Source Dataset Number of Samples Feature Shape Epoch Learning Rate Lambda
-Cls ImageNet 5000 $256\times 256$ 800 1e-4 0.001, 0.0017, 0.003, 0.0035, 0.01
-Seg VOC2012 5000 $256\times 256$ 800 1e-4 0.0005, 0.001, 0.003, 0.007, 0.015
-Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
-CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.15, 10
-TTI Captions from COCO2017 50000 $512\times 512$ 60 1e-4 0.005, 0.01, 0.02, 0.05, 0.2
-Table 7: Training details of the Hyperprior baseline.
-8 Details of baseline methods
-8.1 Packing details
-8.1.1 Semantic segmentation
-In semantic segmentation, an image is pre-processed into multiple $518\times 518$ patches by crop and resize operations.
-These patches are fed into DINOv2, which produces features of shape $1370\times 1536$.
-In the proposed dataset, each image is pre-processed into 2 patches, resulting in features with a shape of $2\times 1370\times 1536$.
-We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
-8.1.2 Depth estimation
-In depth estimation, an image is first horizontally flipped.
-Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2.
-For each image, DINOv2 generates $4\times 1611\times 1536$ features at the split point $SP_{DM}$.
-Thus, the shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$.
-$F_{Dpt}$ is packed in two steps.
-First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width.
-Second, the two $1611\times 6144$ features (original first and then flipped images) are vertically stacked, forming a $3222\times 6144$ feature.
-8.1.3 Text-to-image synthesis
-For text-to-image synthesis, the original feature $F_{TTI}$ has a shape of $16\times 128\times 128$, where 16 represents the number of channels.
-We divide the 16 channels into 4 subgroups.
-Within each subgroup, the $128\times 128$ features are horizontally concatenated to form a $128\times 512$ feature.
-Finally, the four $128\times 512$ features are vertically stacked to produce a final packed feature of size $512\times 512$.
-8.2 Training details
-8.2.1 Training dataset construction
-We construct our training dataset from public datasets.
-The training features are extracted in the same way as the test dataset.
-The source code will be made publicly available, and we encourage researchers to extract features from their own source data.
-As mentioned in the main text, raw features consume a significant amount of storage.
-During training, we encountered storage challenges.
-For example, 20000 original $F_{Dpt}$ features occupy 1474 GB of storage.
-As the demand for additional training data grows, storage requirements could increase substantially, often exceeding the resources available to researchers in academic institutions.
-To address this issue, we crop the features into smaller dimensions to facilitate feature coding training.
-The specific cropping configurations are detailed in Table 7.
-It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training.
-For large models, cropping features to accelerate training is impractical, as downstream models typically require complete semantic information.
-For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation.
-Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
-After feature extraction, the same truncation and quantization operations are applied to the training dataset.
-For the Hyperprior baseline, all features are normalized to the range [0, 1] before training.
-8.2.2 Hyperparameter Setting
-The hyperparameters, such as the number of training epochs, initial learning rate, and $\lambda$ values, are provided in Table 7.
-The number of training epochs is determined based on the size of the training dataset.
-In our experiments, all training processes converge to a stable loss.
-Five different $\lambda$ values are used to achieve various bitrates.
-Task Image Classification Semantic Segmentation Depth Estimation Common Sense Reasoning Text-to-Image Synthesis
-Metric QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec. QP Enc. Dec.
-VTM Baseline 22 267.20 0.23 22 2301.30 1.11 22 14169.36 5.15 22 336.27 0.24 22 107.98 0.16
-27 275.02 0.22 27 2039.34 0.94 27 10862.76 3.54 27 423.84 0.29 27 87.33 0.15
-32 124.40 0.15 32 1399.71 0.52 32 7776.30 2.33 32 362.98 0.24 32 62.11 0.15
-37 28.00 0.13 37 407.30 0.32 37 1590.27 1.14 37 162.05 0.17 37 32.38 0.24
-42 8.55 0.14 42 70.66 0.27 42 348.68 1.17 42 66.28 0.18 42 15.70 0.15
-Metric $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec. $\lambda$ Enc. Dec.
-Hyperprior Baseline 0.01 0.04 0.06 0.015 0.3 0.38 0.12 1.57 2.05 10 0.07 0.12 0.2 0.03 0.03
-0.0035 0.04 0.05 0.007 0.31 0.4 0.05 1.59 2.05 0.07 0.07 0.09 0.05 0.03 0.03
-0.003 0.04 0.05 0.003 0.31 0.39 0.02 1.55 1.88 0.015 0.06 0.08 0.02 0.03 0.03
-0.0017 0.05 0.06 0.001 0.29 0.35 0.005 1.54 1.84 0.0142 0.07 0.08 0.01 0.03 0.03
-0.001 0.05 0.06 0.0005 0.29 0.36 0.001 1.57 1.91 0.01405 0.06 0.08 0.005 0.03 0.03
-Table 8: Encoding and decoding complexity evaluations. “Enc.” and “Dec.” denote encoding time and decoding time, respectively. The running time is measured in seconds.
-$\lambda$-BPFP
-Semantic Segmentation
-Common Sense Reasoning
-Text-to-Image Synthesis
-Figure 6: $\lambda$-BPFP curves and R-A curves of generalizability evaluation on Seg, CSR, and TTI tasks.
-9 Additional experimental results
-$SP_{DM1},QP=42$
-$SP_{DM1},\lambda=0.001$
-$SP_{DM2},QP=42$
-$SP_{DM2},\lambda=0.001$
-$SP_{DM3},QP=42$
-$SP_{DM3},\lambda=0.001$
-$SP_{DM4},QP=42$
-$SP_{DM4},\lambda=0.001$
-$SP_{GS},QP=32$
-$SP_{GS},\lambda=0.01405$
-$SP_{H},QP=42$
-$SP_{H},\lambda=0.005$
-Figure 7: Frequency and CDF visualization of the original truncated and reconstructed features.
-Green and red are for the original truncated features.
-Yellow and black are for the reconstructed features.
-Left: the VTM baseline.
-Right: the Hyperprior baseline.
-9.1 Complexity analysis
-We present the encoding and decoding complexity analyses in Table 8.
-The VTM baseline and Hyperprior baseline are evaluated on an Intel® Xeon® E5-2690 v4 CPU and a single NVIDIA GeForce RTX 4090 GPU, respectively.
-It is important to note that the VTM baseline is a handcrafted codec and cannot be executed on a GPU.
-For the VTM baseline, encoding time increases as QP decreases and feature resolution increases.
-The block-based VTM codec exhaustively examines multiple partitions and modes during encoding, making it significantly more time-consuming than decoding.
-VTM is one of the reference implementations of the VVC standard.
-In practical deployments, a more efficient VVC codec, such as VVEnc, can be used to accelerate encoding.
-For the Hyperprior baseline, the decoding complexity is slightly higher than the encoding complexity.
-Across different bitrates, the encoding and decoding complexities remain comparable.
-9.2 Generalizability analysis
-Here, we visualize the $\lambda$-BPFP curves and the rate-accuracy curves for the cross evaluation to provide a more intuitive presentation of the results.
-The $\lambda$-BPFP curves reveal that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties.
-From the rate-accuracy curves, we observe that coding models trained for one task struggle to achieve comparable performance when applied to another task.
-The generalizability of feature coding methods across multiple tasks has received limited attention in prior research.
-By presenting these results, we aim to highlight this important research direction and encourage further exploration.
-9.3 Ablation on truncation and quantization
-In this subsection, we examine the impact of truncation and quantization on feature coding.
-As shown in Table 10, truncation and quantization individually result in little task accuracy loss.
-When both are applied together, task accuracies experience slight decreases.
-However, applying the VTM baseline to quantized features leads to a significant reduction in accuracy.
-Without truncation, feature values are often limited to a small subset within the full range of [0, 1023].
-In this scenario, during feature coding, the codec’s internal quantization maps values with a small range to a single value, introducing significant distortion.
-By applying truncation, this range is expanded, which helps mitigate distortion caused by VTM’s internal quantization.
-To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37).
-As presented in Table 11, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates.
-Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
-Proposed Features Existing Features
-Source Dataset Split Point Max Min Mean IV GM Split Point Max Min Mean IV GM
-ImageNet $SP_{DS}$ 79.01 -502.04 0.0782 17.04 3.97 $C5$-ReLU 7.81 0 0.4594 9839.49 143.22
-VOC2012 $SP_{DS}$ 94.88 -485.23 0.0780 13.30 4.30 $C5$ 8.00 -4.24 0.2473 7323.72 131.90
-NYUv2 $SP_{DM1}$ 3.25 -2.33 -0.0015 128.91 17.92 $p2$ 10.83 -$1$0.1726 8452.88 159.45
-$SP_{DM2}$ 4.99 -26.08 -0.0005 36.96 10.00 $p3$ 8.35 -9.02 -0.1277 9418.53 168.38
-$SP_{DM3}$ 23.51 -314.26 0.0042 9.61 2.76 $p4$ 9.06 -9.93 -0.0629 8499.73 158.89
-$SP_{DM4}$ 96.04 -493.17 0.0731 12.54 4.35 $p5$ 9.61 -8.22 -0.0295 9402.06 169.13
-Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
-Captions from COCO2017 $SP_{H}$ 3.67 -2.11 0.5664 21130.12 138.53 / / / / / /
-Table 9: Statistics of the proposed features and existing features. (Gradient magnitude values are scaled by $10^{5}$).
-Configuration Cls Seg Dpt CSR
-Truncation Only 100 81.60 0.4992 100
-Quantization Only 99 81.37 0.4982 100
-Truncation + Quantization 100 79.93 0.4941 100
-Quantization + VTM QP22 0 5.35 1.6085 6
-Table 10: Ablations on truncation and quantization operations.
-9.4 Distribution analysis on decoded features
-We visualize the distribution of reconstructed features in Fig.
-7, comparing the two baseline methods at similar bitrates.
-For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range.
-The reconstructed features appear truncated during the feature coding process.
-We attribute this truncation to the Hyperprior’s internal quantization (rounding operation), where the transformed latent variables are directly rounded to integers before entropy coding.
-In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon.
-A similar truncation phenomenon is observed for $SP_{H}$.
-For $SP_{GS}$, the features are encoded at a higher bitrate and the truncation phenomenon is not observed.
-10 Additional dataset analysis
-In addition to the visualizations provided in the main text, we present further statistical analyses of the proposed dataset.
-10.1 Distribution analysis
-We compare the statistical characteristics of the proposed features with those of existing features by randomly sampling 10 examples from each dataset and reporting their maximum, minimum, and mean values in Table 9.
-Overall, the proposed features exhibit a broader distribution range compared to existing features.
-In particular, for multi-scale features, the distribution range expands in deeper network layers, whereas existing features exhibit smaller variations.
-Moreover, although the proposed features tend to be more asymmetric, their mean values remain closer to zero.
-Image Classification Semantic Segmentation
-[-5, 5] [-30, 30] [-5, 5] [-30, 30]
-BPFE Acc. BPFE Acc. BPFE mIoU BPFE mIoU
-2.90 100 0.38 91 2.84 79.83 0.37 76.38
-2.05 100 0.06 24 1.98 79.40 0.08 63.46
-1.15 100 0.02 14 1.10 78.68 0.01 39.32
-Depth Estimation Common Sense Reasoning
-[-5, 5], [-5, 5] [-20, 20], [-30, 30] [-2, 2] [-10, 10]
-BPFE RMSE BPFE RMSE BPFE Acc. BPFE Acc.
-1.42 0.6490 0.58 0.8569 3.16 98 0.84 98
-0.73 0.9604 0.17 1.1360 2.30 99 0.15 82
-0.33 1.2453 0.04 1.3073 1.44 99 0.04 20
-Table 11: Ablation results on truncation regions. $SP_{DM1}$ and $SP_{DM2}$ features are truncated into [-1, 1] and [-2, 2] in all cases, respectively.
-10.2 Redundancy analysis
-To evaluate spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), where gradients are derived using the Sobel operator.
-Before these computations, both pixel and feature values are quantized to 10-bit integers.
-The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range from 7323.72 to 9839.49.
-A similar trend is observed in gradient magnitude.
-These broader variations underscore the higher diversity of the proposed features.
-Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features.
-For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude.
-This observation highlights the necessity and importance of incorporating textual features in the dataset.
-10.3 Visualization
-We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig.
-8 and Fig.
-9.
-These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer.
-To avoid redundancy in the main paper, these visualizations have been moved here.
-$SP_{DS},F_{Cls}$
-$SP_{DS},F_{Seg}$
-Figure 8: Visualization of the frequency distribution and CDF for the $F_{Cls}$ and $F_{Seg}$ features.
-$SP_{DS},F_{Cls}$
-$SP_{DS},F_{Seg}$
-Figure 9: Visualization of the original feature blocks and their corresponding DCT blocks for the $F_{Cls}$ and $F_{Seg}$ features.